### v1

# Multiscale scattered data analysis in samplet coordinates

Abstract.
We study multiscale scattered data interpolation schemes for globally supported radial basis functions, with a focus on the Matérn class. The multiscale approximation is constructed through a sequence of residual corrections, where radial basis functions with different lengthscale parameters are employed to capture varying levels of detail. To apply this approach to large data sets, we suggest to represent the resulting generalized Vandermonde matrices in samplet coordinates. Samplets are localized, discrete signed measures exhibiting vanishing moments and allow for the sparse approximation of generalized Vandermonde matrices issuing from a vast class of radial basis functions. Given a quasi-uniform set of $N$ data sites, and local approximation spaces with geometrically decreasing dimension, the full multiscale system can be assembled with cost $\mathcal{O}(N\log N)$. We prove that the condition numbers of the linear systems at each level remain bounded independent of the particular level, allowing us to use an iterative solver with a bounded number of iterations for the numerical solution. Hence, the overall cost of the proposed approach is $\mathcal{O}(N\log N)$. The theoretical findings are accompanied by extensive numerical studies in two and three spatial dimensions.
1. Introduction
Interpolation using radial basis functions (RBFs) is a widely recognized technique for fitting functions based on scattered data in $\mathbb{R}^{d}$, see [10, 13, 14] for some surveys. RBF interpolation allows for the construction of approximation spaces in arbitrary dimensions and with varying levels of smoothness. These spaces are well known to provide high-quality approximations, see [11, 25]. However, the actual computation of an RBF interpolant usually entails a significant computational cost and the corresponding linear system is typically ill-conditioned. Even worse, as the convergence order increases, the associated problems become successively ill-conditioned. Moreover, using globally supported RBFs in interpolation leads to densely populated matrices, making the problem computationally infeasible for larger datasets. To address the latter, techniques such as interpolation using compactly supported RBFs have been developed [19, 24]. Unfortunately, even interpolation with these kernels becomes either computationally too expensive, if unscaled kernels are used, or do not even lead to convergence, if the support is scaled proportional to the fill-distance. Motivated by solving this trade-off principle, the works [3, 5, 9, 21], among many others, leverage the idea of a multiscale method where approximations on different scales are computed in an residual correction scheme. To date, the most general convergence theorems for multiscale interpolation using positive definite, compactly supported RBFs have been developed in [21]. The advantage of the multiscale method is that the condition numbers of the resulting linear systems at each level remain independent of the particular level. Thus achieving linear convergence in the number of used levels with relatively low cost.
In this article, we study the RBF interpolation problem using globally supported RBFs. We address the challenge of dealing with densely-populated interpolation matrices by employing samplet compression, a technique introduced in [7]. Samplets are discrete signed measures constructed such that all polynomials up to a certain degree vanish. This vanishing moment property allows for the compression of generalized Vandermonde matrices issuing from asymptotically smooth RBFs, i.e., they behave like smooth functions apart from zero. Hence, In samplet coordinates, the resulting interpolation matrices become quasi-sparse, which means that they can be compressed such that only a sparse matrix remains. Furthermore, samplets allow for a respective sparse matrix algebra, see [8], and provide a meaningful interpretation of sparsity constraints for scattered data, see [2]. However, as the samplet transform of a given generalized Vandermonde matrix is an isometry up to the compression error, it does not alleviate the ill-conditioning. Therefore, to mitigate issues related to ill-conditioning, we adopt a multiscale interpolation approach that exploits scaled versions of the same RBF, which, in our case, remains globally supported. Hence, recovering the results from [21] for the class of Matérn kernels.
The remainder of this article is organized as follows. In Section 2, we revisit the classical multiscale interpolation algorithm, detailing the relevant subspaces and the class of radial basis functions considered in this work. In Section 3 we carry over a known result for compactly supported functions to globally supported RBFs: The condition number of the kernel matrix can be bounded uniformly in the number of sites, if we apply the appropriate scaling. Section 4 provides a brief introduction of samplets, with a focus on the key properties of the basis transformation and compression. In Section 5, we present the multiscale interpolation algorithm within the samplet basis and show an example of the resulting generalized Vandermonde matrix pattern. Finally, in Section 6, we provide numerical tests in both two and three spatial dimensions. Concluding remarks are stated in Section 7.
2. Classical multiscale method
Let $\Omega\subseteq\mathbb{R}^{d}$ be a domain and $X=\{{\boldsymbol{x}}_{1},\ldots,{\boldsymbol{x}}_{N}\}\subset\Omega$ be a set of data sites with cardinality $N=\#X$. Moreover, let $f\in C(\Omega)$ be the target function generating the data $f_{1},\dots,f_{N}$. Associated to the set of sites $X$ there are two measures, the fill-distance of $X$ in $\Omega$ and the separation radius of $X$,
$h_{X,\Omega}$ $\mathrel{\mathrel{\mathop{:}}=}\sup_{{\boldsymbol{x}}\in\Omega}\min_{{\boldsymbol{x}}\in X}\|{\boldsymbol{x}}-{\boldsymbol{x}}_{j}\|_{2}$
$q_{X}$ $\mathrel{\mathrel{\mathop{:}}=}\frac{1}{2}\min_{j\neq k}\|{\boldsymbol{x}}_{j}-{\boldsymbol{x}}_{k}\|_{2}.$
We call $X$ quasi-uniform if there is a constant $c_{\text{qu}}>0$ such that
$q_{X}\leq h_{X,\Omega}\leq c_{\text{qu}}q_{X}.$
Our goal is to interpolate or approximate the target function by means of radial basis functions (RBF) using only the given data $({\boldsymbol{x}}_{1},f_{1}),\ldots,({\boldsymbol{x}}_{N},f_{N})$. A function $\Phi\colon\mathbb{R}^{d}\to\mathbb{R}$ is said to be radial, if and only if there exists a univariate function $\phi\colon[0,\infty)\to\mathbb{R}$ such that $\Phi({\boldsymbol{x}})=\phi(\|{\boldsymbol{x}}\|_{2})$. The radial function $\Phi$ is strictly positive definite, if and only if there holds for any choice of distinct points ${\boldsymbol{\xi}}_{1},\ldots,{\boldsymbol{\xi}}_{n}$, $n\in\mathbb{N}$, that
${\boldsymbol{\alpha}}^{\intercal}[\Phi({\boldsymbol{\xi}}_{i}-{\boldsymbol{\xi}}_{j})]_{i,j=1}^{n}{\boldsymbol{\alpha}}>0\quad\text{for all }{\boldsymbol{\alpha}}\in\mathbb{R}^{n}\setminus\{{\boldsymbol{0}}\}.$
In this case, the kernel
$K({\boldsymbol{x}_{i}},{\boldsymbol{x}_{j}})\mathrel{\mathrel{\mathop{:}}=}\Phi({\boldsymbol{x}}_{i}-{\boldsymbol{x}}_{j})$
is the reproducing kernel of a uniquely determined reproducing kernel Hilbert space, its native space, that we denote by $\big{(}\mathcal{N}_{\Phi},\langle\cdot,\cdot\rangle_{\mathcal{N}_{\Phi}}\big{)}$. In this context, we will use the function $\Phi$ and its associated kernel $K$ synonymously. A popular class of strictly positive RBFs are the Matérn kernels or Sobolev splines $\Phi_{\nu}\colon\mathbb{R}^{d}\to\mathbb{R}$, dependent on a hyper-parameter $\nu>d/2$. These kernels are defined via the radial function
(1) $\phi_{\nu}(r)=\frac{2^{1-\nu}}{\Gamma(\nu)}r^{\nu-\frac{d}{2}}K_{\nu-\frac{d}{2}}(r),\quad r\geq 0,$
where $\Gamma$ is the Riemann Gamma function and $K_{\beta}$ is the modified Bessel function of the second kind, see [12] for example.
It is well-known that the Fourier transform of this radial function behaves like
(2) $\widehat{\Phi}(\boldsymbol{\omega})=(1+\|\boldsymbol{\omega}\|_{2}^{2})^{-\nu-\frac{d}{2}},\quad\boldsymbol{\omega}\in\mathbb{R}^{d},$
and that, in turn, the native space of the corresponding Matérn kernel is norm-equivalent to the classical Sobolev space $H^{\nu+\frac{d}{2}}(\mathbb{R}^{d})$, see [20, Theorem 6.13] for example.
In this instationary setting it is known that the interpolant
$s\in\mathcal{S}\mathrel{\mathrel{\mathop{:}}=}\operatorname{span}\{\Phi(\cdot-{\boldsymbol{x}}):{\boldsymbol{x}}\in X\}\subset\mathcal{N}_{\Phi}$
converges towards the target function $f$ as $h_{X,\Omega}\to 0$. However, this improved approximation comes with higher and higher numerical cost for the iterative solution due to the ill-conditioning. To alleviate this cost, we can rescale the kernel, i.e., we introduce the parameter $\delta>0$ and define
(3) $\Phi_{\delta}\mathrel{\mathrel{\mathop{:}}=}\delta^{-d}\Phi\bigg{(}\frac{\cdot}{\delta}\bigg{)}.$
In the stationary setting we couple $\delta$ proportional to $h_{X,\Omega}$, i.e., there is a coupling parameter $\eta>0$ such that $\delta=\eta h_{X,\Omega}$. In this case we can control the numerical costs to compute the interpolant but cannot expect convergence. This is known as the trade-off principle.
The idea of the multiscale method is now to introduce a hierarchy
(4) $X_{1},X_{2},\ldots,X_{L}=X,$
where we refer to the subscript of $X_{\ell}$ as level with cardinalities $N_{\ell}$ and fill-distances $h_{\ell}\mathrel{\mathrel{\mathop{:}}=}h_{X_{\ell},\Omega}$. Although in many applications the sets will be nested, for the multiscale method itself, this is not necessary. We assume that the cardinalities of the sets $X_{\ell}$ are increasing with the level, i.e., $N_{\ell}\leq N_{\ell+1}$ or, in other words, $h_{\ell}\geq h_{\ell+1}$. We quantify this relation in Theorem 2.1. Now, we use the stationary setting on each level $\ell$. We have level-dependent $\delta_{\ell}>0$ and level-dependently rescaled kernel $K_{\ell}$ given by
$K_{\ell}(\boldsymbol{x},\boldsymbol{y})\mathrel{\mathrel{\mathop{:}}=}\Phi_{\ell}(\boldsymbol{x}-\boldsymbol{y})=\delta_{\ell}^{-d}\Phi\left(\frac{\boldsymbol{x}-\boldsymbol{y}}{\delta_{\ell}}\right),\quad\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d}.$
Each set $X_{\ell}$ in the multiscale hierarchy (4), together with the respective kernel, gives rise to a subspace
(5) $\mathcal{S}_{\ell}\mathrel{\mathrel{\mathop{:}}=}\operatorname{span}\{\Phi_{\ell}(\cdot-{\boldsymbol{x}}):{\boldsymbol{x}}\in X_{\ell}\}\subset\mathcal{N}_{\Phi_{\ell}},$
also called local approximation space.
With the multiscale sequence of points and the associated kernels, we can decompose the original approximation problem into $L$ steps using a telescoping sum
$\left.s_{1}\right|_{X_{1}}\mathrel{=\mathrel{\mathop{:}}}\left.w_{1}\right|_{X_{1}}$ $=\left.f\right|_{X_{1}},$
$\left.w_{2}\right|_{X_{2}}$ $=\left.\left(f-w_{1}\right)\right|_{X_{2}},$
[…]
$w_{L}|_{X_{L}}$ $=\bigg{(}f-\sum_{\ell=1}^{L-1}w_{\ell}\bigg{)}\bigg{|}_{X_{L}}.$
The intuition is that at the $\ell$-th step we compute the approximation to the error function or residual of level $\ell-1$, i.e.,
(6) $f-(w_{1}+\cdots+w_{l-1})$
in the space $\mathcal{S}_{\ell}$, given as
(7) $w_{\ell}({\boldsymbol{x}})=\sum_{i=1}^{N_{\ell}}c_{\ell,i}\Phi_{\ell}({\boldsymbol{x}}-{\boldsymbol{x}}_{i}),\quad\boldsymbol{x}\in\Omega,$
with a coefficient vector ${\boldsymbol{c}}_{\ell}\in\mathbb{R}^{N_{\ell}}$.
From this, we obtain the multiscale decomposition
$s_{L}\mathrel{\mathrel{\mathop{:}}=}(w_{1}+\cdots+w_{L})|_{X}=f|_{X},$
which means that $s_{L}\in\cup_{\ell=1}^{L}\mathcal{S}_{\ell}$ matches $f$ at each of the data sites in $X$.
In practice, the multiscale approximation process begins with a coarse set of points, where the approximation problem (7) is solved. Next, the error function (6) is computed, and finally, the latter is approximated at a finer scale. Interpolating the residual at level $\ell$ can be written in the following form:
$\boldsymbol{K}_{\ell,\ell}\boldsymbol{c}_{\ell}=\boldsymbol{f}_{\ell}-\sum_{\ell^{\prime}=1}^{\ell-1}\boldsymbol{K}_{\ell,\ell^{\prime}}\boldsymbol{c}_{\ell^{\prime}},$
where
$\boldsymbol{K}_{\ell,\ell^{\prime}}\mathrel{\mathrel{\mathop{:}}=}\big{[}\Phi_{\ell^{\prime}}\big{(}{\boldsymbol{x}}_{i}^{(\ell)}-\boldsymbol{x}_{j}^{(\ell^{\prime})}\big{)}\big{]}_{i,j},\quad\ell,\ell^{\prime}=1,\ldots,L.$
Therefore, the algorithm to solve the multiscale interpolation consists of solving the following triangular system, whose entries are matrices of different sizes:
(8) $\left[\begin{array}[]{ccccc}\boldsymbol{K}_{1,1}&&&&\\ \boldsymbol{K}_{2,1}&\boldsymbol{K}_{2,2}&&&\\ \boldsymbol{K}_{3,1}&\boldsymbol{K}_{3,2}&\boldsymbol{K}_{3,3}&&\\ \vdots&\vdots&\cdots&\ddots&\\ \boldsymbol{K}_{L,1}&\boldsymbol{K}_{L,2}&\cdots&\boldsymbol{K}_{L,L-1}&\boldsymbol{K}_{L,L}\end{array}\right]\left[\begin{array}[]{c}\boldsymbol{c}_{1}\\ \boldsymbol{c}_{2}\\ \boldsymbol{c}_{3}\\ \vdots\\ \boldsymbol{c}_{L}\end{array}\right]=\left[\begin{array}[]{c}\boldsymbol{f}_{1}\\ \boldsymbol{f}_{2}\\ \boldsymbol{f}_{3}\\ \vdots\\ \boldsymbol{f}_{L}\end{array}\right]$
The resulting matrix is in general a densely-populated populated lower triangular matrix as a result of the choice of globally supported RBFs for the approximation at each level. The samplet compression plays a crucial role here, rendering the matrix quasi-sparse. This technique is introduced in Section 4, and the linear system is reformulated in the samplet basis in Section 5.
There holds the following convergence result, see, e.g., [21, 22].
Theorem 2.1.
Let $\Omega\subseteq\mathbb{R}^{d}$ be a bounded domain with Lipschitz boundary. Let $X_{1},X_{2},\dots$ be a sequence of point sets in $\Omega$ with fill-distances $h_{1},h_{2},\dots$ satisfying
(9) $c\mu h_{\ell}\leq h_{\ell+1}\leq\mu h_{\ell}$
for $\ell=1,2,\dots$ with fixed $\mu\in(0,1)$, $c\in(0,1]$ and $h_{1}$ sufficiently small. Let $\Phi\colon\mathbb{R}^{d}\to\mathbb{R}$ be a reproducing kernel for $H^{s}(\mathbb{R}^{d})$, i.e., its Fourier transform satisfies
$c_{1}(1+\|{\boldsymbol{\omega}}\|_{2}^{2})^{-s}\leq\widehat{\Phi}({\boldsymbol{\omega}})\leq c_{2}(1+\|{\boldsymbol{\omega}}\|_{2}^{2})^{-s},\quad{\boldsymbol{\omega}}\in\mathbb{R}^{d},$
and let $\Phi_{\ell}$ be the rescaled kernel as in (3) with scale factor $\delta_{\ell}=\eta h_{\ell}$, $\eta>0$. Let $f\in H^{s}(\Omega)$. Then there exist constants $C$ and $C_{1}>0$ such that
$\|f-s_{L}\|_{L_{2}(\Omega)}\leq C(C_{1}\mu^{s}+C_{1}\eta^{-s})^{L}\|f\|_{H^{s}(\Omega)}.$
3. Condition number bound for generalized Vandermonde matrices
We now show that the condition number $\operatorname{cond}_{2}$ of the generalized Vandermonde matrix $\boldsymbol{K}_{\ell,\ell}$ of the scaled kernel can be bounded by a constant, independently of the level $\ell$. A similar result in the context of the multiscale method for compactly supported RBFs has been derived in [18] using techniques of [15, 16]. We will reuse these techniques, with appropriate modifications. Furthermore, for a lighter notation, we will drop the subscripts $\ell$ in this section.
It goes without saying that, such a bound can not be established for general sets of sites $X$. Hence, for this section, we assume that $X$ is quasi-uniform and the parameters of the multilevel method are chosen as in Theorem 2.1.
We remark that the algebraic decay of the Fourier transform of the Matérn kernel (2) immediately yields for the scaled RBF $\Phi_{\delta}$ that
(10) $\widehat{\Phi_{\delta}}(\boldsymbol{\omega})=(1+\delta^{2}\|\boldsymbol{\omega}\|_{2}^{2})^{-\nu-\frac{d}{2}},\quad\boldsymbol{\omega}\in\mathbb{R}^{d}.$
We will bound the smallest and largest eigenvalue of the matrix ${\boldsymbol{K}}$ separately. We follow the ideas of [18, 15, 16].
Theorem 3.1.
Let $\Omega\subseteq\mathbb{R}^{d}$ be a bounded domain and $X\subseteq\Omega$ be a quasi-uniform set of sites with fill-distance $h_{X,\Omega}$ and separation radius $q_{X}$. Let $\Phi$ be the Matérn kernel and Let $\Phi_{\delta}=\delta^{-d}\Phi(\cdot/\delta)$ be the rescaled kernel where $\delta=\eta h_{X,\Omega}$ with overlap parameter $\eta>1$. Then there exists a constant $C=C(\Phi,d,s)$ such that the bound for the smallest eigenvalue $\lambda_{min}$ of $\boldsymbol{K}=[\Phi_{\delta}(\boldsymbol{x}_{i}-\boldsymbol{x}_{j})]_{i,j}$
$\lambda_{min}({\boldsymbol{K}})\geq C(\eta c_{\textnormal{qu}})^{d-2s}\delta^{-d}$
holds. The largest eigenvalue of ${\boldsymbol{K}}$ can be bounded from above by
$\lambda_{max}({\boldsymbol{K}})\leq\widetilde{C}\delta^{-d},$
with a constant $\widetilde{C}=C(\Phi,d,\nu,\eta,c_{\textnormal{qu}})$.
Proof.
We set
$\widetilde{\boldsymbol{K}}=\delta^{d}{\boldsymbol{K}}=\bigg{[}\Phi\bigg{(}\frac{{\boldsymbol{x}}_{i}-{\boldsymbol{x}}_{j}}{\delta}\bigg{)}\bigg{]}_{i,j}.$
The lower bound follows immediately from [18, Proof of Proposition 3.4].
We only need to prove the upper bound for $\lambda_{max}({\boldsymbol{K}})$. We use Gershgorin’s circle theorem, see, e.g. [23, Theorem 5.1] for $\widetilde{\boldsymbol{K}}$. Since the diagonal elements in this matrix are $\Phi(\boldsymbol{0})$, independent of the line $j$, we can, without loss of generality, set $j=1$ and $\boldsymbol{x}_{1}=\boldsymbol{0}$. We have
(11) $|\lambda-\Phi(\boldsymbol{0})|$ $\leq\sum_{k=2}^{N}\bigg{|}\Phi\bigg{(}\frac{\boldsymbol{0}-\boldsymbol{x}_{k}}{\delta}\bigg{)}\bigg{|}=\sum_{k=2}^{N}\bigg{|}\Phi\bigg{(}\frac{\boldsymbol{x}_{k}}{\delta}\bigg{)}\bigg{|}.$
To bound the right hand side of this estimate, we introduce the annuli $E_{n}$ given by
$E_{n}=\{\boldsymbol{x}\in\mathbb{R}^{d}\;:\;nq_{X}\leq\|\boldsymbol{x}\|_{2}<(n+1)q_{X}\}.$
Each $\boldsymbol{x}_{k}$ is contained in exactly one $E_{n}$ and we can estimate
#$\#\{\boldsymbol{x}_{k}\in E_{n}\}\leq 3^{d}n^{d-1},$
see, e.g., [20, Proof of Theorem 12.3]. Inserting this into (11) yields
$|\lambda_{max}(\widetilde{\boldsymbol{K}})-\Phi(\boldsymbol{0})|$ $\leq\sum_{k=2}^{N}\bigg{|}\Phi\bigg{(}\frac{\boldsymbol{x}_{k}}{\delta}\bigg{)}\bigg{|}$
$\leq\sum_{n=1}^{\infty}\sum_{\boldsymbol{x}_{k}\in E_{n}}\bigg{|}\Phi\bigg{(}\frac{\boldsymbol{x}_{k}}{\delta}\bigg{)}\bigg{|}$
(12) $\leq\sum_{n=1}^{\infty}\#\{\boldsymbol{x}_{k}\in E_{n}\}\sup_{\boldsymbol{x}\in E_{n}}\bigg{|}\Phi\bigg{(}\frac{\boldsymbol{x}}{\delta}\bigg{)}\bigg{|}.$
Next, we bound $\sup_{\boldsymbol{x}\in E_{n}}|\Phi({\boldsymbol{x}}/{\delta})|$. We use the specific form of $\Phi$, cp. (1), and the fact that the function $r\mapsto r^{\beta}K_{\beta}(r)$ is non-increasing on $(0,\infty)$ for any $\beta\in\mathbb{R}$, see [20, Corollary 5.12]. This yields
$\frac{\Gamma(\nu)}{2^{1-\nu}}\sup_{\boldsymbol{x}\in E_{n}}\bigg{|}\Phi\bigg{(}\frac{\boldsymbol{x}}{\delta}\bigg{)}\bigg{|}$ $=\sup_{\boldsymbol{x}\in E_{n}}\bigg{(}\frac{\|\boldsymbol{x}\|_{2}}{\delta}\bigg{)}^{\nu-\frac{d}{2}}K_{\nu-\frac{d}{2}}\bigg{(}\frac{\|\boldsymbol{x}\|_{2}}{\delta}\bigg{)}$
$\leq\bigg{(}\frac{nq_{X}}{\delta}\bigg{)}^{\nu-\frac{d}{2}}K_{\nu-\frac{d}{2}}\bigg{(}\frac{nq_{X}}{\delta}\bigg{)}$
$\leq\bigg{(}\frac{nq_{X}}{\delta}\bigg{)}^{\nu-\frac{d}{2}}\sqrt{2\pi}\bigg{(}\frac{\delta}{nq_{X}}\bigg{)}^{\frac{1}{2}}\exp\bigg{(}-\frac{nq_{X}}{\delta}\bigg{)}\exp\Bigg{(}\bigg{(}\nu-\frac{d}{2}\bigg{)}^{2}\frac{2\delta}{2nq_{X}}\Bigg{)},$
where we used the asymptotic behavior [20, Lemma 5.13] of the modified Bessel function to arrive at the last inequality. Inserting this into (3), using the quasi-uniformity of $X$ and the coupling of $\delta$ to $h_{X,\Omega}$, as in Theorem 2.1, yields
$|\lambda_{max}(\widetilde{\boldsymbol{K}})-\Phi(\boldsymbol{0})|$ $\leq\sum_{n=1}^{\infty}\#\{\boldsymbol{x}_{k}\in E_{n}\}\sup_{\boldsymbol{x}\in E_{n}}\bigg{|}\Phi\bigg{(}\frac{\boldsymbol{x}}{\delta}\bigg{)}\bigg{|}$
$\leq\sum_{n=1}^{\infty}3^{d}n^{d-1}\sqrt{2\pi}\eta^{\frac{d}{2}-\nu}(\eta c_{\text{qu}})^{\frac{1}{2}}n^{\nu-\frac{d}{2}-\frac{1}{2}}\exp\bigg{(}-\frac{n}{\eta}+\frac{(\nu-\frac{d}{2})^{2}\eta c_{q}u}{2n}\bigg{)}$
$\leq\sqrt{2\pi}3^{d}\eta^{\frac{d}{2}-\nu}(\eta c_{\text{qu}})^{\frac{1}{2}}\sum_{n=1}^{\infty}n^{\nu-\frac{d}{2}-\frac{3}{2}}\exp\bigg{(}-\frac{n}{\eta}+\frac{(\nu-\frac{d}{2})^{2}\eta c_{q}u}{2n}\bigg{)}.$
Since the series on the right hand side of the estimate converges, this yields the claimed bound for $\lambda_{max}$. ∎
The derived bounds on $\lambda_{min}$ and $\lambda_{max}$ lead to the boundedness of the condition number of the generalized Vandermonde matrix for every level, uniformly in the number of data sites and the scaling parameter.
Corollary 3.2.
For every $\ell\in\mathbb{N}$ the condition number $\operatorname{cond}_{2}(\boldsymbol{K}_{\ell,\ell})$ is bounded independently of $N_{\ell}$ and $\delta_{\ell}$.
4. Samplets for scattered data approximation
Samplets are discrete signed measures, which exhibit vanishing moments. We briefly recall their underlying concepts as introduced in [7]. To this end, we consider the image of the spaces $\mathcal{S}_{\ell}$, $\ell=1,\ldots,L$, under the Riesz isomorphism $\mathcal{J}\colon\mathcal{N}_{\Phi_{\ell}}\to\mathcal{N}^{\prime}_{\Phi_{\ell}}$ and define
(13) $\mathcal{S}_{\ell}^{\prime}\mathrel{\mathrel{\mathop{:}}=}\operatorname{span}\{\mathcal{J}\Phi_{\ell}(\cdot-{\boldsymbol{x}}):{\boldsymbol{x}}\in X_{\ell}\}=\operatorname{span}\{\delta_{\boldsymbol{x}}:{\boldsymbol{x}}\in X_{\ell}\},$
by the identity $\mathcal{J}\Phi_{\ell}(\cdot-{\boldsymbol{x}})=\delta_{\boldsymbol{x}}$, where $\delta_{\boldsymbol{x}}$ is the point evaluation functional at ${\boldsymbol{x}}\in\Omega$. On the spaces $\mathcal{S}_{\ell}^{\prime}$, we introduce an inner product different from the canonical one by the defining property $\langle\delta_{{\boldsymbol{x}}_{i}},\delta_{{\boldsymbol{x}}_{j}}\rangle_{\mathcal{S}_{\ell}^{\prime}}\mathrel{\mathrel{\mathop{:}}=}\delta_{ij}$ for ${\boldsymbol{x}}_{i},{\boldsymbol{x}}_{j}\in X_{\ell}$.
For the sake of a more lightweight notation, we drop the subscript $\ell$ of the spaces $\mathcal{S}_{\ell}^{\prime}$ for the remainder of this paragraph and remark that each space $\mathcal{S}_{\ell}^{\prime}$ will in general have a different multiresolution analysis, where the maximum level $J$ depends on $\ell$. Given a multiresolution analysis
(14) $\mathcal{V}_{0}^{\prime}\subset\mathcal{V}_{1}^{\prime}\subset\cdots\subset\mathcal{V}_{J}^{\prime}=\mathcal{S}^{\prime},$
we keep track of the increment of information between two consecutive levels $j$ and $j+1$. Since there holds $\mathcal{V}_{j}^{\prime}\subset\mathcal{V}_{j+1}^{\prime}$, we may orthogonally decompose
$\mathcal{V}_{j+1}^{\prime}=\mathcal{V}_{j}^{\prime}\overset{\perp}{\oplus}\mathcal{W}_{j}^{\prime}$
for a certain detail space $\mathcal{W}_{j}^{\prime}$. In analogy to wavelet nomenclature, we call the elements of a basis of $\mathcal{V}_{0}^{\prime}$ scaling distributions and the elements of a basis of one of the spaces $\mathcal{W}_{j}^{\prime}$ samplets. This name is motivated by the idea that the basis distributions in $\mathcal{W}_{j}^{\prime}$ are supported at a small subsample or samplet of the data sites in $X$. The collection of the bases of $\mathcal{W}_{j}^{\prime}$ for $j=0,\ldots,J-1$ together with a basis of $\mathcal{V}_{0}^{\prime}$ is called a samplet basis for $\mathcal{S}^{\prime}$. A samplet basis can be constructed such that it exhibits vanishing moments of order $q+1$, i.e.,
(15) $(p,\sigma_{j,k})_{\Omega}=0\ \text{for all}\ p\in\mathcal{P}_{q},$
where $\sigma_{j,k}\in\mathcal{W}_{j}^{\prime}$ is a samplet and $\mathcal{P}_{q}\mathrel{\mathrel{\mathop{:}}=}\operatorname{span}\{{\boldsymbol{x}}^{\boldsymbol{\alpha}}:\|\boldsymbol{\alpha}\|_{1}\leq q\}$ denotes the space of polynomials of total degree at most $q$.
The samplet construction for $\mathcal{S}^{\prime}$ is based on a cluster tree for the set of data sites $X$, i.e., a tree $\mathcal{C}$ with root $X$ such that each node $\tau\in\mathcal{C}$ is the disjoint union of its children. Such a cluster tree $\mathcal{C}$ directly induces a support based hierarchical clustering of the subspace $\mathcal{S}^{\prime}$, spanned by the Dirac-$\delta$-distributions supported at the data sites in $X$. With a slight abuse of notation, we will refer to this cluster tree also by $\mathcal{C}$ and to its elements by $\tau$. Given a cluster tree $\mathcal{C}$, a samplet basis for $\mathcal{S}^{\prime}$, where $\dim\mathcal{S}^{\prime}=N$, can be constructed with cost $\mathcal{O}(N)$. Assuming furthermore, that $\mathcal{C}$ is a balanced binary tree with maximum level $J=\lceil\log_{2}(N)\rceil$, the samplet basis exhibits the following properties, see [7] but also [6, 17].
Theorem 4.1.
The samplet basis $\bigcup_{j=0}^{J}\{\sigma_{j,k}\}_{k}$ forms an orthonormal basis in $\mathcal{S}^{\prime}$, satisfying the following properties:
  1. (1)
    There holds $c^{-1}2^{j}\leq\operatorname{dim}\mathcal{W}_{j}^{\prime}\leq c2^{j}$ for a constant $1<c\leq 2$.
  2. (2)
    The samplets have vanishing moments of order $q+1$, i.e., $(p,\sigma_{j,k})_{\Omega}=0$ for all $p\in\mathcal{P}_{q}(\Omega)$, where $\mathcal{P}_{q}(\Omega)$ is the space of polynomials up to degree $q$.
  3. (3)
    The coefficient vector ${\boldsymbol{\omega}}_{j,k}=\big{[}\omega_{j,k,i}\big{]}_{i}$ of the samplet $\sigma_{j,k}$ satisfies the bound
    $\|{\boldsymbol{\omega}}_{j,k}\|_{1}\leq c2^{(J-j)/2}$
    for a constant $20<c<$ 2.
  4. (4)
    Given $f\in C^{q+1}(O)$, for some open set $O\supset\operatorname{supp}\sigma_{j,k}$, there holds
    $|\sigma_{j,k}(f)|\leq\bigg{(}\frac{d}{2}\bigg{)}^{q+1}\frac{(\operatorname{diam}\operatorname{supp}\sigma_{j,k})^{q+1}}{(q+1)!}\|f\|_{C^{q+1}(O)}\|{\boldsymbol{\omega}}_{j,k}\|_{1}.$
  5. (5)
    For quasi-uniform sets $X$, there holds $\operatorname{diam}\operatorname{supp}\sigma_{j,k}\leq c2^{-j/d}$ for some constant $c>0$.
A direct consequence of the orthonormality of the samplet basis is that the samplet transform
$[\sigma_{j,k}]_{j,k}={\boldsymbol{T}}[\delta_{{\boldsymbol{x}}_{i}}]_{i=1}^{N}$
satisfies ${\boldsymbol{T}}^{\intercal}{\boldsymbol{T}}={\boldsymbol{T}}{\boldsymbol{T}}^{\intercal}={\boldsymbol{I}}\in\mathbb{R}^{N\times N}$. If the samplet transform ${\boldsymbol{T}}$ is computed recursively adhering the hierarchy induced by the cluster tree $\mathcal{C}$, its cost is of order $\mathcal{O}(N)$.
Employing the Riesz isometry, the samplet basis gives rise to a basis for $\mathcal{S}\subset\mathcal{N}_{\Phi}$. To this end, consider a samplet
$\sigma_{j,k}=\sum_{i=1}^{N}\omega_{j,k,i}\delta_{{\boldsymbol{x}}_{i}}.$
The samplet can be identified with the function
$\psi_{j,k}\mathrel{\mathrel{\mathop{:}}=}\sum_{i=1}^{N}\omega_{j,k,i}\Phi({\boldsymbol{x}}_{i}-\cdot)\in\mathcal{S}$
by means of the Riesz isometry. The vanishing moment property (15) then translates to
$\langle\psi_{j,k},f\rangle_{\mathcal{N}_{\Phi}}=0$
for any $f\in\mathcal{N}_{\Phi}$ which satisfies $f|_{O}\in\mathcal{P}_{q}$ for any open and convex set $O$ with $\operatorname{supp}(\sigma_{j,k})\subset O\cap\Omega$.
Furthermore, there holds
(16) $\big{[}\big{\langle}\psi_{j,k},\psi_{j^{\prime},k^{\prime}}\big{\rangle}_{\mathcal{N}_{\Phi}}\big{]}_{j,j^{\prime},k,k^{\prime}}={\boldsymbol{T}}{\boldsymbol{K}}{\boldsymbol{T}}^{\intercal}\mathrel{=\mathrel{\mathop{:}}}{\boldsymbol{K}}^{\Sigma}.$
This means that the Gramian of the embedded samplet basis in $\mathcal{S}$ is just the samplet transformed generalized Vandermonde matrix. For asymptotically smooth kernels $K$, i.e. that there exist $C,r>0$ such that for all $({\boldsymbol{x}},{\boldsymbol{y}})\in(\Omega\times\Omega)\setminus\Delta$
(17) $\bigg{|}\frac{\partial^{|\boldsymbol{\alpha}|+|\boldsymbol{\beta}|}}{\partial{\boldsymbol{x}}^{\boldsymbol{\alpha}}\partial{\boldsymbol{y}}^{\boldsymbol{\beta}}}K({\boldsymbol{x}},{\boldsymbol{y}})\bigg{|}\leq C\frac{(|\boldsymbol{\alpha}|+|\boldsymbol{\beta}|)!}{r^{|\boldsymbol{\alpha}|+|\boldsymbol{\beta}|}\|{\boldsymbol{x}}-{\boldsymbol{y}}\|_{2}^{|\boldsymbol{\alpha}|+|\boldsymbol{\beta}|}}$
uniformly in $\boldsymbol{\alpha},\boldsymbol{\beta}\in\mathbb{N}^{d}$ apart from the diagonal $\Delta\mathrel{\mathrel{\mathop{:}}=}\{({\boldsymbol{x}},{\boldsymbol{y}})\in\Omega\times\Omega:{\boldsymbol{x}}={\boldsymbol{y}}\}$, the matrix ${\boldsymbol{K}}^{\Sigma}$ becomes quasi-sparse and can efficiently be computed. More precisely, there holds the following result, see [7].
Theorem 4.2.
Let $X$ be quasi-uniform with #$\#X=N$ and set all coefficients of the generalized Vandermonde matrix ${\boldsymbol{K}}^{\Sigma}$ from (16) to zero which satisfy the admissibility condition
(18) $\operatorname{dist}(\tau,\tau^{\prime})\geq\rho\max\{\operatorname{diam}(\tau),\operatorname{diam}(\tau^{\prime})\},\quad\rho>0,$
where $\tau$ is the cluster supporting $\sigma_{j,k}$ and $\tau^{\prime}$ is the cluster supporting $\sigma_{j^{\prime},k^{\prime}}$, respectively. Then, there exists a constant $C>0$, such that
(19) $\frac{\big{\|}{\boldsymbol{K}}^{\Sigma}-{\boldsymbol{K}}^{\Sigma}_{\varepsilon}\big{\|}_{F}}{\big{\|}{\boldsymbol{K}}^{\Sigma}\big{\|}_{F}}\leq Cm_{q}(r\rho/d)^{-2(q+1)},$
where $m_{q}=\binom{q+d}{d}$ is the dimension of $\mathcal{P}_{q}$ , and the compressed matrix has $\mathcal{O}(m_{q}^{2}N\log N)$ nonzero coefficients.
5. Multiscale algorithm in samplet coordinates
In this section, we transform the interpolation problem from the original basis, referred to as the natural basis, to the samplet basis, which consists of the scaling distributions and the samplets, as introduced in Section 4. In what follows, we denote the samplet transform for $\mathcal{S}^{\prime}_{\ell}$ by ${\boldsymbol{T}}_{\ell}\in\mathbb{R}^{N_{\ell}\times N_{\ell}}$ and recall that ${\boldsymbol{T}}^{\intercal}_{\ell}{\boldsymbol{T}}_{\ell}={\boldsymbol{T}}_{\ell}{\boldsymbol{T}}^{\intercal}_{\ell}={\boldsymbol{I}}\in\mathbb{R}^{N_{\ell}\times N_{\ell}}$. Defining the block transform
(20) $\mathcal{T}\mathrel{\mathrel{\mathop{:}}=}\left[\begin{array}[]{cccc}{\boldsymbol{T}}_{1}&&&\\ &{\boldsymbol{T}}_{2}&&\\ &&\ddots&\\ &&&{\boldsymbol{T}}_{L}\end{array}\right],$
there holds $\mathcal{T}^{\intercal}\mathcal{T}=\mathcal{T}\mathcal{T}^{\intercal}={\boldsymbol{I}}$. Hence, the linear system (8) is equivalent to
$\mathcal{T}^{\intercal}\mathcal{T}\left[\begin{array}[]{cccc}\boldsymbol{K}_{1,1}&&&\\ \boldsymbol{K}_{2,1}&\boldsymbol{K}_{2,2}&&\\ \vdots&\cdots&\ddots&\\ \boldsymbol{K}_{L,1}&\cdots&\boldsymbol{K}_{L,L-1}&\boldsymbol{K}_{L,L}\end{array}\right]\mathcal{T}^{\intercal}\mathcal{T}\left[\begin{array}[]{c}\boldsymbol{c}_{1}\\ \boldsymbol{c}_{2}\\ \boldsymbol{c}_{3}\\ \vdots\\ \boldsymbol{c}_{L}\end{array}\right]=\mathcal{T}^{\intercal}\mathcal{T}\left[\begin{array}[]{c}\boldsymbol{f}_{1}\\ \boldsymbol{f}_{2}\\ \boldsymbol{f}_{3}\\ \vdots\\ \boldsymbol{f}_{L}\end{array}\right]$
Letting $\boldsymbol{K}^{\Sigma}_{\ell,\ell^{\prime}}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{T}}_{\ell}\boldsymbol{K}_{\ell,\ell^{\prime}}{\boldsymbol{T}}^{\intercal}_{\ell^{\prime}}$, $\boldsymbol{c}^{\Sigma}_{\ell}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{T}}_{\ell}\boldsymbol{c}_{\ell}$ and $\boldsymbol{f}^{\Sigma}_{\ell}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{T}}_{\ell}\boldsymbol{f}_{\ell}$ and dropping the inverse block transform, we arrive at the samplet transformed system
(21) $\left[\begin{array}[]{ccccc}\boldsymbol{K}^{\Sigma}_{1,1}&&&&\\ \boldsymbol{K}^{\Sigma}_{2,1}&\boldsymbol{K}^{\Sigma}_{2,2}&&&\\ \boldsymbol{K}^{\Sigma,3}_{1}&\boldsymbol{K}^{\Sigma}_{3,2}&\boldsymbol{K}^{\Sigma}_{3,3}&&\\ \vdots&\vdots&\cdots&\ddots&\\ \boldsymbol{K}^{\Sigma}_{L,1}&\boldsymbol{K}^{\Sigma}_{L,2}&\cdots&\boldsymbol{K}^{\Sigma}_{L,L-1}&\boldsymbol{K}^{\Sigma}_{L,L}\end{array}\right]\left[\begin{array}[]{c}\boldsymbol{c}^{\Sigma}_{1}\\ \boldsymbol{c}^{\Sigma}_{2}\\ \boldsymbol{c}^{\Sigma}_{3}\\ \vdots\\ \boldsymbol{c}^{\Sigma}_{L}\end{array}\right]=\left[\begin{array}[]{c}\boldsymbol{f}^{\Sigma}_{1}\\ \boldsymbol{f}^{\Sigma}_{2}\\ \boldsymbol{f}^{\Sigma}_{3}\\ \vdots\\ \boldsymbol{f}^{\Sigma}_{L}\end{array}\right],$
which is equivalent to (8). In particular, since $\mathcal{T}$ is an isometry, the condition number of the system (21) is identical to that of (8).
Furthermore, we remark that, for a geometric sequence $\{N_{\ell}\}_{\ell}$, the cost for applying $\mathcal{T}$ and $\mathcal{T}^{\intercal}$, respectively, remains linear in $N_{L}$. Each block ${\boldsymbol{K}}^{\Sigma}_{\ell,\ell^{\prime}}$, $\ell\leq\ell^{\prime}$, can be computed with cost $\mathcal{O}(N_{\ell^{\prime}}+N_{\ell}\log N_{\ell})$ if $\{N_{\ell}\}_{\ell}$ forms a geometric sequence, cp. [1, Proof of Theorem 3.16]. Moreover, there are $\mathcal{O}(\log N_{\ell^{\prime}})$ blocks per row, resulting in a computational cost of $\mathcal{O}(N_{\ell^{\prime}}\log N_{\ell^{\prime}})$ per row by the geometric sequence property of $\{N_{\ell}\}_{\ell}$. Summing up the cost for each row, exploiting the geometric sequence again, the overall computation cost for the assembly of the linear system (21) is bounded by $\mathcal{O}(N_{L}\log N_{L})$. To give an idea of the sparsity pattern of the samplet compressed system (21), Figure 1 visualizes the matrix pattern $\boldsymbol{K}^{\Sigma}_{\ell,\ell^{\prime}}$ using three levels of points $X_{1},X_{2},X_{3}$ arranged on a regular grid for the unit square, with cardinalities of 4 225, 66 049, and 1 050 625, respectively.
Figure 1. Example of the sparsity patterns of the matrices $\boldsymbol{K}^{\Sigma}_{\ell,\ell^{\prime}}$ for ,$\ell,\ell^{\prime}\in\{1,2,3\}$ in the linear system (21).
6. Numerical results
We present numerical examples of the proposed multiscale interpolation in samplets coordinates in both two and three dimensional settings. The parameters for the samplet compression are $\rho$, $q$, and $\kappa$. The parameter $\rho$ corresponds to the admissibility condition (18), which steers the a-priorily truncated matrix blocks. A larger $\rho$ results in a larger distance between clusters before they are admissible for compression. In our analysis, we adopt $\rho=d$, so that the bound in Equation (19) is exponentially governed by the number of vanishing moments. This number of vanishing moments corresponds to $q+1$, as specified in (15). The higher $q$, the more polynomial exactness is gained. Lastly, $\kappa$ serves as an a-posteriori compression threshold. After applying the samplet transform to the generalized Vandermonde matrix, we truncate entries smaller than $\kappa$, resulting in a sparser matrix representation.
To measure the approximation quality, we consider the relative interpolation errors
$\text{error}_{p,{\ell}}=\frac{\|f-s_{\ell}\|_{p}}{\|f\|_{p}}$
for $p=2$ and $p=\infty$ at each of the levels $\ell=1,\ldots,L$. Furthermore, we compute the order of convergence, given by
$\text{order}_{p,{\ell}}=\frac{\log{\left(\displaystyle\frac{\text{error}_{p,{\ell+1}}}{{\text{error}_{p,{\ell}}}}\right)}}{\log{\left(\displaystyle\frac{h_{\ell+1}}{h_{\ell}}\right)}}.$
We investigate three different scenarios in our numerical tests. The first two are adapted from [21]: a smooth function on the unit square and a non-smooth function on the L-shaped domain. Finally, we extend to a three dimensional setting issuing from the Lucy surface model of the Stanford 3D Scanning Repository [^1] https://graphics.stanford.edu/data/3Dscanrep/.
All computations have been performed at the Centro Svizzero di Calcolo Scientifico (CSCS) on a single node of the Alps cluster [^2] https://www.cscs.ch/computers/alps with two AMD EPYC 7742 @2.25GHz CPUs, up to 420GB of main memory and up to 20 cores. The implementation of the samplet compression is available as software package FMCA [^3] https://github.com/muchip/fmca.
6.1. Franke’s function in two spatial dimensions
We interpolate Franke’s function, as introduced in [4], defined by
(22) $f(x,y)=$ $\frac{3}{4}\exp\left(-\frac{(9x-2)^{2}}{4}-\frac{(9y-2)^{2}}{4}\right)+\frac{3}{4}\exp\left(-\frac{(9x+1)^{2}}{49}-\frac{(9y+1)}{10}\right)$
$+\frac{1}{2}\exp\left(-\frac{(9x-7)^{2}}{4}-\frac{(9y-3)^{2}}{4}\right)-\frac{1}{5}\exp\left(-(9x-4)^{2}-(9y-7)^{2}\right)$
on the unit square $\Omega=[0,1]^{2}$. To this end, we generate 10 sets of nested points within this domain, distributing them on a regular grid with size $h_{j}=2^{-j}$. The number of points per level and the corresponding fill-distances are provided in Table 1.
$\ell$ $N_{\ell}$ $h$
1 9 0.5
2 25 0.25
3 81 0.125
4 289 0.0625
5 1 089 0.03125
6 4 225 0.015625
7 16 641 0.0078125
8 66 049 0.00390625
9 263 169 0.00195312
10 1 050 625 0.000976562
Table 1. Number of points per level and the corresponding fill-distances on $\Omega=[0,1]^{2}$ for the interpolation of Franke’s function.
We have conducted computations using a-posteriori compression thresholds of $\kappa=10^{-4},10^{-6},10^{-8}$. For $\kappa=10^{-4}$, the error remains larger than $10^{-6}$, due to an increased compression error. Setting $\kappa=10^{-6}$, we have achieved an error of the order of $10^{-8}$. Further reducing the threshold to $\kappa=10^{-8}$ results in an error of $10^{-9}$. Even so, the latter results in a significantly worse compression rate of the generalized Vandermonde matrices.
Therefore, we detail the results using $\kappa=10^{-6}$ for the threshold. We approximate the solution with the ${C}^{1}$-smooth Matérn-3/2 RBF and the ${C}^{2}$-smooth Matérn-5/2 RBF and different values of $\gamma$. The error is evaluated at a fine grid with mesh-size $h=2^{-11}$. The accuracy of the conjugate gradient method is set to $10^{-6}$. Increasing the accuracy does not result in smaller errors but increases the number of iterations. Interpolation error, order and compression rate (%nz) as well as (non-preconditioned) conjugate gradient iterations (CG) are presented in the Tables 25.
Matérn-3/2 RBF
$\gamma=0.25$ $\gamma=0.5$
$\ell$ $\text{error}_{2}$ $\text{order}_{2}$nz CG $\text{error}_{2}$ $\text{order}_{2}$nz CG
1 $5.67\cdot 10^{-1}$ 55 5 $4.84\cdot 10^{-1}$ 55 6
2 $1.38\cdot 10^{-1}$ 2.04 36 8 $5.29\cdot 10^{-2}$ 2.19 36 13
3 $3.16\cdot 10^{-2}$ 2.12 43 9 $9.20\cdot 10^{-3}$ 2.53 43 25
4 $6.76\cdot 10^{-3}$ 2.23 47 10 $1.15\cdot 10^{-3}$ 3.00 47 34
5 $1.48\cdot 10^{-3}$ 2.19 22 10 $2.51\cdot 10^{-4}$ 2.53 22 39
6 $3.45\cdot 10^{-4}$ 2.10 7.8 10 $6.08\cdot 10^{-5}$ 2.05 7.6 39
7 $8.30\cdot 10^{-5}$ 2.06 2.3 10 $1.30\cdot 10^{-5}$ 2.22 2.3 39
8 $1.46\cdot 10^{-5}$ 2.51 0.62 10 $1.14\cdot 10^{-6}$ 3.51 0.62 39
9 $3.14\cdot 10^{-6}$ 2.22 0.16 10 $1.04\cdot 10^{-7}$ 3.45 0.16 39
10 $7.89\cdot 10^{-7}$ 1.99 0.04 10 $4.47\cdot 10^{-8}$ 1.21 0.04 39
Table 2. Results for the interpolation of Franke’s function using the ${C}^{1}$-smooth Matérn-3/2 RBF and $\rho=2$, $q+1=5$, $\kappa=10^{-6}$ as samplet compression parameters.
Matérn-3/2 RBF
$\gamma=0.75$ $\gamma=1$
$\ell$ $\text{error}_{2}$ $\text{order}_{2}$nz CG $\text{error}_{2}$ $\text{order}_{2}$nz CG
1 $4.84\cdot 10^{-1}$ 55 6 $4.88\cdot 10^{-1}$ 55 6
2 $4.67\cdot 10^{-2}$ 3.37 36 17 $4.51\cdot 10^{-2}$ 3.43 36 20
3 $7.97\cdot 10^{-3}$ 2.55 43 41 $7.72\cdot 10^{-3}$ 2.55 43 48
4 $5.70\cdot 10^{-4}$ 3.81 47 64 $3.60\cdot 10^{-4}$ 4.43 47 97
5 $1.09\cdot 10^{-4}$ 2.38 23 90 $5.75\cdot 10^{-5}$ 3.31 25 156
6 $2.38\cdot 10^{-5}$ 2.19 8.2 96 $1.16\cdot 10^{-5}$ 2.31 8.9 187
7 $4.57\cdot 10^{-6}$ 2.38 2.5 98 $2.07\cdot 10^{-6}$ 2.82 2.7 195
8 $4.77\cdot 10^{-7}$ 3.26 0.68 99 $2.27\cdot 10^{-7}$ 3.19 0.74 197
9 $5.97\cdot 10^{-8}$ 3.00 0.18 99 $7.85\cdot 10^{-8}$ 1.53 0.19 199
10 $4.63\cdot 10^{-8}$ 0.37 0.05 99 $7.30\cdot 10^{-8}$ 0.10 0.05 200
Table 3. Results for the interpolation of Franke’s function using the ${C}^{1}$-smooth Matérn-3/2 RBF and $\rho=2$, $q+1=5$, $\kappa=10^{-6}$ as samplet compression parameters.
Matérn-5/2 RBF
$\gamma=0.25$ $\gamma=0.5$
$\ell$ $\text{error}_{2}$ $\text{order}_{2}$nz CG $\text{error}_{2}$ $\text{order}_{2}$nz CG
1 $5.45\cdot 10^{-1}$ 55 5 $4.72\cdot 10^{-1}$ 55 6
2 $1.12\cdot 10^{-1}$ 2.28 36 8 $4.51\cdot 10^{-2}$ 1.39 36 16
3 $2.27\cdot 10^{-2}$ 2.14 43 9 $7.73\cdot 10^{-3}$ 2.54 43 32
4 $4.54\cdot 10^{-3}$ 2.32 47 10 $6.92\cdot 10^{-4}$ 3.48 47 42
5 $1.00\cdot 10^{-3}$ 2.18 22 10 $1.49\cdot 10^{-4}$ 2.88 22 48
6 $2.49\cdot 10^{-4}$ 2.01 7.6 10 $3.42\cdot 10^{-5}$ 2.12 7.7 48
7 $6.13\cdot 10^{-5}$ 2.02 2.3 10 $6.88\cdot 10^{-6}$ 2.32 2.3 48
8 $6.06\cdot 10^{-6}$ 3.34 0.62 9 $8.31\cdot 10^{-7}$ 3.05 0.62 48
9 $1.15\cdot 10^{-6}$ 2.40 0.16 9 $6.15\cdot 10^{-8}$ 3.76 0.16 48
10 $3.07\cdot 10^{-7}$ 1.91 0.04 9 $1.57\cdot 10^{-8}$ 1.97 0.04 48
Table 4. Results for the interpolation of Franke’s function using the ${C}^{2}$-smooth Matérn-5/2 RBF and $\rho=2$, $q+1=5$, $\kappa=10^{-6}$ as samplet compression parameters.
Matérn-5/2 RBF
$\gamma=0.75$ $\gamma=1$
$\ell$ $\text{error}_{2}$ $\text{order}_{2}$nz CG $\text{error}_{2}$ $\text{order}_{2}$nz CG
1 $4.79\cdot 10^{-1}$ 55 6 $9.16\cdot 10^{-1}$ 55 7
2 $4.41\cdot 10^{-2}$ 3.44 36 20 $4.69\cdot 10^{-2}$ 4.29 36 25
3 $7.13\cdot 10^{-3}$ 2.63 43 57 $7.47\cdot 10^{-3}$ 2.65 43 82
4 $3.29\cdot 10^{-4}$ 4.43 47 97 $1.63\cdot 10^{-4}$ 2.19 47 172
5 $6.14\cdot 10^{-5}$ 2.42 23 139 $3.96\cdot 10^{-5}$ 2.04 24 298
6 $1.22\cdot 10^{-5}$ 2.33 8.0 151 $1.34\cdot 10^{-5}$ 1.56 8.8 370
7 $2.11\cdot 10^{-6}$ 2.53 2.4 157 $2.06\cdot 10^{-6}$ 2.70 2.7 402
8 $2.57\cdot 10^{-7}$ 3.03 0.70 159 $2.39\cdot 10^{-7}$ 3.10 0.73 416
9 $4.06\cdot 10^{-8}$ 2.66 0.17 161 $7.56\cdot 10^{-8}$ 1.66 0.19 430
10 $3.04\cdot 10^{-8}$ 0.42 0.04 162 $7.01\cdot 10^{-8}$ 0.11 0.05 430
Table 5. Results for the interpolation of Franke’s function using the ${C}^{2}$-smooth Matérn-5/2 RBF and $\rho=2$, $q+1=5$, $\kappa=10^{-6}$ as samplet compression parameters.
The tables clearly demonstrate the expected exponential decay of the approximation error with increasing level for both error measures and both RBFs under consideration. Furthermore, we see the effect of a larger value of $\gamma$, which to a greater lengthscale parameter of the RBF, resulting in a larger condition number. For increasing $\gamma$, we require successively more iterations for the conjugate gradient method to achieve the prescribed accuracy, with up to 430 iterations for $\gamma=1$ and the Matérn-5/2 RBF. The deterioration of the convergence for $\gamma=0.75,1$ on the last level is caused by the samplet compression error, which is a directly to the chosen compression threshold $\kappa$.
Figure 2 shows a visualization of the solution (left) and the residuals (right) at levels $\ell=1,4,7,10$ for $\gamma=1$ using the Matérn-3/2 RBF.
Refer to captionRefer to caption$\ell=1$Refer to captionRefer to caption$\ell=4$Refer to captionRefer to caption$\ell=7$Refer to captionRefer to caption$\ell=10$
Figure 2. Solution (left) and corresponding residuals (right) for the interpolation of Franke’s function for levels $\ell=1,4,7,10$.
6.2. Non-smooth function
We consider the L-shaped domain $\Omega=\left[-\frac{1}{2},\frac{1}{2}\right]^{2}\setminus\left(0,\frac{1}{2}\right]^{2}$ where we introduce polar coordinates $x=r\cos\phi$, $y=r\sin\phi$, with $r\geq 0$ and $\phi\in[\pi/2,2\pi]$. In this test, we aim at interpolating the function
$u(r,\phi)=-r^{\frac{2}{3}}\sin\left(\frac{2\phi-\pi}{3}\right),$
which solves the Laplace equation $\Delta u=0$. We select $h_{j}=2^{-j}$ as before. Given the reduced smoothness of the function, we choose the ${C}^{0}$-smooth Matérn-1/2 RBF for its approximation. The error is evaluated at a fine grid with mesh-size $h=2^{-9}$ and the accuracy of the conjugate gradient method is set to $10^{-6}$.
We have conducted experiments using $\kappa=10^{-4},10^{-6},10^{-8}$. While the magnitude of the errors remained consistent across these values, the compression rate in the final level increased as $\kappa$ decreased. Specifically for $\kappa=10^{-4}$, the compression rate is around $0.13\%$, for $\kappa=10^{-6}$, it is about $0.30\%$ and for $\kappa=10^{-8}$ it is approximately $0.43\%$. Consequently, we chose a higher threshold to reduce computational cost and storage. The number of points per level and the corresponding fill-distances are provided in Table 6.
$\ell$ $N_{\ell}$ $h$
1 21 0.25
2 65 0.125
3 225 0.0625
4 833 0.03125
5 3 201 0.015625
6 12 545 0.0078125
7 49 665 0.00390625
8 197 633 0.00195312
Table 6. Number of points per level and the corresponding fill-distances on the L-shaped domain for the interpolation of the solution to the Laplace equation.
The results for $\gamma=0.5$ and $\gamma=1$ are summarized in Table 7.
$\gamma=0.5$
$\ell$ $\text{error}_{2}$ $\text{error}_{\infty}$ $\text{order}_{2}$ $\text{order}_{\infty}$nz CG
1 $4.37\cdot 10^{-2}$ $5.98\cdot 10^{-2}$ 52 9
2 $1.38\cdot 10^{-2}$ $3.17\cdot 10^{-2}$ 1.67 0.92 49 21
3 $5.33\cdot 10^{-3}$ $2.04\cdot 10^{-2}$ 1.37 0.64 38 28
4 $1.83\cdot 10^{-3}$ $1.29\cdot 10^{-2}$ 1.54 0.66 18 34
5 $5.88\cdot 10^{-4}$ $8.15\cdot 10^{-3}$ 1.64 0.67 6.2 36
6 $1.85\cdot 10^{-4}$ $5.04\cdot 10^{-3}$ 1.67 0.69 1.8 37
7 $5.87\cdot 10^{-5}$ $3.11\cdot 10^{-3}$ 1.66 0.70 0.51 37
8 $1.89\cdot 10^{-5}$ $1.61\cdot 10^{-3}$ 1.63 0.95 0.13 37
$\gamma=1$
$\ell$ $\text{error}_{2}$ $\text{error}_{\infty}$ $\text{order}_{2}$ $\text{order}_{\infty}$nz CG
1 $3.39\cdot 10^{-2}$ $6.06\cdot 10^{-2}$ 48 13
2 $1.27\cdot 10^{-2}$ $3.94\cdot 10^{-2}$ 1.41 0.62 49 28
3 $4.15\cdot 10^{-3}$ $2.50\cdot 10^{-2}$ 1.62 0.66 37 49
4 $1.29\cdot 10^{-3}$ $1.58\cdot 10^{-2}$ 1.68 0.66 18 73
5 $3.97\cdot 10^{-4}$ $9.94\cdot 10^{-3}$ 1.70 0.67 6.0 93
6 $1.23\cdot 10^{-4}$ $6.13\cdot 10^{-3}$ 1.70 0.70 1.8 102
7 $3.83\cdot 10^{-5}$ $3.72\cdot 10^{-3}$ 1.68 0.72 0.12 104
8 $1.20\cdot 10^{-5}$ $1.95\cdot 10^{-3}$ 1.68 0.76 0.09 105
Table 7. Results for the interpolation of the solution of the Laplace equation on the L-shaped domain using the ${C}^{0}$-smooth Matérn-1/2 RBF and $\rho=2$, $q+1=5$, $\kappa=10^{-4}$ as samplet compression parameters.
The compression rates are similar to the previous example, also in case of the less smooth Matérn-1/2 RBF. The approximation error and the corresponding order are reduced, however. The number of conjugate gradient iterations stays rather low with a maximum of 37 iterations for $\gamma=0.5$ and of 105 iterations for $\gamma=1$. This can be explained by the slower decay of the Matérn-1/2 kernel’s eigenvalues, resulting in less ill-conditioned generalized Vandermonde matrices.
Figure 3 illustrates the solution (left) and the corresponding residuals (right) at levels $\ell=2,4,6,8$ with $\gamma=0.5$.
Refer to captionRefer to caption$\ell=2$Refer to captionRefer to caption$\ell=4$Refer to captionRefer to caption$\ell=6$Refer to captionRefer to caption$\ell=8$
Figure 3. Solution (left) and corresponding residuals (right) for the interpolation of the solution to the Laplace equation on the L-shaped domain for levels $\ell=2,4,6,8$ using $\gamma=1$ and $\kappa=10^{-4}$.
6.3. Function on a point cloud in three spatial dimensions
As final example, we consider a function defined on a uniform resampling of the Lucy surface model of the Stanford 3D Scanning Repository. To construct the local approximation spaces, we selected random subsamples of the resulting point cloud, ensuring a specific cardinality for each set. We start with an initial set of 20 points, and at each subsequent level, the cardinality is increased by a factor of 16. In this study, we present results for both nested and non-nested point sets. For the nested case, the indices of the points are randomly shuffled, and subsamples are obtained by selecting the first $n$ points, where $n$ represents the desired cardinality. In contrast, for the non-nested case, each subsample is randomly selected from the entire dataset independently, i.e., without enforcing the inclusion of points from previous levels. Unlike previous tests, this experiment is more general, as the points are not arranged on a uniform grid, they are not necessarily nested, and they are not necessarily quasi-uniform. As test function, we e interpolate
$f({\boldsymbol{x}})=\frac{1}{\sqrt[4]{\|{\boldsymbol{x}}-{\boldsymbol{x}}_{0}\|_{2}}+10^{-4}}$
The point ${\boldsymbol{x}}_{0}$ is selected near the torch that Lucy holds in her left hand. To optimize performance, we use a relatively small value for the parameter $\gamma$, reduce the number of vanishing moments, and set a higher a-posteriori threshold. These adjustments help decrease both computational time and the number of non-zero entries. However, a very small $\gamma$ can compromise the precision, while a very large $\gamma$ increases the number of conjugate gradient iterations, thereby increasing computational time. A good trade-off is given by $\gamma=0.2$. Furthermore, we employed a diagonally preconditioned conjugate gradient method with accuracy $10^{-6}$ to solve the linear system, which reduces the number of iterations by a factor of about three compared to the unpreconditioned case. The number of points per level and the corresponding fill-distances are provided in Table 8.
$\ell$ $N_{\ell}$ $h$ (nested) $h$ (not-nested)
1 20 0.25 0.34
2 320 0.15 0.13
3 5 120 0.061 0.064
4 81 920 0.017 0.018
5 1 310 720 0.0051 0.0050
Table 8. Number of points per level and corresponding fill-distances for the interpolation on the Lucy point cloud.
The error is evaluated on a finer set of 2 000 000 points, which are randomly selected once from the original dataset. The evaluation set remains fixed throughout all experiments to ensure consistency in error measurement. The results using the Matérn-1/2 and Matérn-3/2 RBFs are shown in Tables 9 and 10.
$\gamma=0.2$
nested not-nested
$\ell$ $\text{error}_{2}$ $\text{order}_{2}$nz CG $\text{error}_{2}$ $\text{order}_{2}$nz CG
1 $5.98\cdot 10^{-1}$ 50 8 $4.49\cdot 10^{-1}$ 50 9
2 $1.60\cdot 10^{-1}$ 2.58 44 26 $1.27\cdot 10^{-1}$ 1.31 42 28
3 $3.08\cdot 10^{-2}$ 1.83 6.5 49 $2.25\cdot 10^{-2}$ 2.44 6.9 45
4 $5.22\cdot 10^{-3}$ 1.39 0.51 56 $2.36\cdot 10^{-3}$ 1.78 0.50 65
5 $7.94\cdot 10^{-4}$ 1.56 0.03 57 $2.51\cdot 10^{-4}$ 1.74 0.03 69
Table 9. Results for the interpolation of a test function on the Lucy point cloud for the $C^{0}$-smooth Matérn-1/2 RBF using $\rho=3$, $q+1=3$, $\kappa=10^{-4}$ as samplet compression parameters.
$\gamma=0.2$
nested not-nested
$\ell$ $\text{error}_{2}$ $\text{order}_{2}$nz CG $\text{error}_{2}$ $\text{order}_{2}$nz CG
1 $6.16\cdot 10^{-1}$ 50 11 $4.50\cdot 10^{-1}$ 50 12
2 $1.48\cdot 10^{-1}$ 2.79 45 74 $1.14\cdot 10^{-1}$ 1.43 44 79
3 $2.92\cdot 10^{-2}$ 1.80 6.5 214 $1.73\cdot 10^{-2}$ 2.66 7.0 247
4 $5.41\cdot 10^{-3}$ 1.32 0.51 341 $1.32\cdot 10^{-3}$ 2.03 0.51 398
5 $8.57\cdot 10^{-4}$ 1.53 0.03 349 $1.61\cdot 10^{-4}$ 1.64 0.03 430
Table 10. Results for the interpolation of a test function on the Lucy point cloud for the $C^{1}$-smooth Matérn-3/2 RBF using $\rho=3$, $q+1=3$, $\kappa=10^{-4}$ as samplet compression parameters.
We observe that the $C^{0}$-smooth Matérn-1/2 RBF performs better in this scenario, primarily due to the improved conditioning of the associated generalized Vandermonde matrix. In both cases, the approximation error reaches the compression error of $10^{-4}$.
A visualization of the solution on all five levels $\ell=1,\ldots,5$ using the Matérn-1/2 RBF and not-nested points is depicted in Figure 4.
Refer to captionRefer to caption$\ell=1$$\ell=2$Refer to captionRefer to caption$\ell=3$$\ell=4$Refer to caption$\ell=5$
Figure 4. Interpolant of the test function on the Lucy point cloud for levels $\ell=1,\ldots,5$ using the Matérn-1/2 RBF.
7. Conclusion
We have proposed an efficient multiscale approach for scattered data interpolation with globally supported RBFs in combination with the samplet matrix compression. In particular, we have extended existing results on the boundedness of the condition numbers of the generalized Vandermonde matrices within the multiscale approximation towards globally supported RBFs. For appropriately chosen lengthscale parameters, the condition numbers stay independent of the level, which allows for an efficient solution of the resulting linear system by the conjugate gradient method with a bounded number of iterations for a given accuracy. The resulting algorithm has an overall cost of $\mathcal{O}(N\log N)$ for $N$ quasi-uniform data sites, given that the points per level form a geometric sequence. The numerical results demonstrate that the presented approach achieves a similar accuracy to the multiscale approach by compactly supported RBFs, while overcoming previous limitations that made the use of globally supported RBFs impractical for large-scale problems. We have demonstrated that the present multiscale approach can handle datasets with millions of data in both two and three spatial dimensions. As test cases, we have considered smooth and non-smooth data in two dimensions and a complex geometry in three spatial dimensions. In all cases, we achieved accurate results that would have been computationally infeasible without the samplet matrix compression.
Acknowledgement. SA and MM were funded by the SNSF starting grant “Multiresolution methods for unstructured data” (TMSGI2_211684).

### v2

@@
 Abstract.
-We study multiscale scattered data interpolation schemes for globally supported radial basis functions, with a focus on the Matérn class.
-The multiscale approximation is constructed through a sequence of residual corrections, where radial basis functions with different lengthscale parameters are employed to capture varying levels of detail.
-To apply this approach to large data sets, we suggest to represent the resulting generalized Vandermonde matrices in samplet coordinates.
+We study multiscale scattered data interpolation schemes for globally supported radial basis functions with focus on the Matérn class.
+The multiscale approximation is constructed through a sequence of residual corrections, where radial basis functions with different lengthscale parameters are combined to capture varying levels of detail.
+We prove that the condition numbers of the the diagonal blocks of the corresponding multiscale system remain bounded independently of the particular level, allowing us to use an iterative solver with a bounded number of iterations for the numerical solution.
+Employing an appropriate diagonal scaling, the multiscale system becomes well conditioned.
+We exploit this fact to derive a general error estimate bounding the consistency error issuing from a numerical approximation of the multiscale system.
+To apply the multiscale approach to large data sets, we suggest to represent each level of the multiscale system in samplet coordinates.
 Samplets are localized, discrete signed measures exhibiting vanishing moments and allow for the sparse approximation of generalized Vandermonde matrices issuing from a vast class of radial basis functions.
-Given a quasi-uniform set of $N$ data sites, and local approximation spaces with geometrically decreasing dimension, the full multiscale system can be assembled with cost $\mathcal{O}(N\log N)$.
-We prove that the condition numbers of the linear systems at each level remain bounded independent of the particular level, allowing us to use an iterative solver with a bounded number of iterations for the numerical solution.
-Hence, the overall cost of the proposed approach is $\mathcal{O}(N\log N)$.
+Given a quasi-uniform set of $N$ data sites, and local approximation spaces with exponentially decreasing dimension, the samplet compressed multiscale system can be assembled with cost $\mathcal{O}(N\log^{2}N)$.
+The overall cost of the proposed approach is $\mathcal{O}(N\log^{2}N)$.
 The theoretical findings are accompanied by extensive numerical studies in two and three spatial dimensions.
@@
 These spaces are well known to provide high-quality approximations, see [18, 37].
-However, the actual computation of an RBF interpolant usually entails a significant computational cost and the corresponding linear system is typically ill-conditioned.
+Even so, the computation of an RBF interpolant usually entails a significant computational cost and the corresponding linear system is typically ill-conditioned.
 Even worse, as the convergence order increases, the associated problems become successively ill-conditioned.
@@
 To address the latter, techniques such as interpolation using compactly supported RBFs have been developed [27, 36].
-Unfortunately, even interpolation with these kernels becomes either computationally too expensive, if unscaled kernels are used, or do not even lead to convergence, if the support is scaled proportional to the fill-distance.
-Motivated by solving this trade-off principle, the works [3, 5, 9, 21], among many others, leverage the idea of a multiscale method where approximations on different scales are computed in an residual correction scheme.
+Unfortunately, even interpolation with these RBFs either becomes computationally too expensive, if unscaled kernels are used, or does not even lead to convergence, if the support is scaled proportional to the fill-distance.
+Motivated by balancing this trade-off principle, the works [4, 6, 15, 29], among many others, leverage the idea of a multiscale method where approximations on different scales are computed in a residual correction scheme.
 To date, the most general convergence theorems for multiscale interpolation using positive definite, compactly supported RBFs have been developed in [29].
 The advantage of the multiscale method is that the condition numbers of the resulting linear systems at each level remain independent of the particular level.
-Thus achieving linear convergence in the number of used levels with relatively low cost.
+An effective approach to address the computational challenges of large datasets is the greedy kernel method [33].
+The greedy kernel method systematically selects a representative subset of the data to construct a sparse approximation.
+An alternative strategy involves deep kernel methods [3, 23, 34, 32].
+These methods leverage deep neural networks to learn adaptive feature representations, thereby constructing kernels that are tuned to the intrinsic structure of medium to high-dimensional data.
+Further, kernel multigrid methods [35, 13, 7] have been developed for the efficient numerical solution of partial differential equations on complex domains such as manifolds or point cloud surfaces.
+These techniques use specialized kernels, such as polyharmonic or homogeneous kernels, along with localized Lagrange bases.
 In this article, we study the RBF interpolation problem using globally supported RBFs.
@@
 Furthermore, samplets allow for a respective sparse matrix algebra, see [11], and provide a meaningful interpretation of sparsity constraints for scattered data, see [2].
+Unlike approaches that rely on the specific kernel, samplets can be applied to a broad class of positive definite kernels.
 However, as the samplet transform of a given generalized Vandermonde matrix is an isometry up to the compression error, it does not alleviate the ill-conditioning.
 Therefore, to mitigate issues related to ill-conditioning, we adopt a multiscale interpolation approach that exploits scaled versions of the same RBF, which, in our case, remains globally supported.
-Hence, recovering the results from [21] for the class of Matérn kernels.
+Especially, we extend the results from [29] for compactly supported RBFs to the class of Matérn kernels.
+As a consequence, the diagonally scaled multiscale system becomes well conditioned.
+We exploit this fact to derive a general error estimate bounding the consistency error issuing from a numerical approximation of the multiscale system.
+The theoretical results are corroborated by extensive numerical studies in two and three spatial dimensions.
 The remainder of this article is organized as follows.
 In Section 2, we revisit the classical multiscale interpolation algorithm, detailing the relevant subspaces and the class of radial basis functions considered in this work.
-In Section 3 we carry over a known result for compactly supported functions to globally supported RBFs: The condition number of the kernel matrix can be bounded uniformly in the number of sites, if we apply the appropriate scaling.
+In Section 3 we carry over a known result for compactly supported functions to globally supported RBFs: The condition number of the generalized Vandermonde matrix can be bounded uniformly in the number of sites, if an appropriate scaling is applied.
+Section 4 derives a general approximation error estimate for the case that the block matrix in the multiscale system is numerically approximated.
 Section 5 provides a brief introduction of samplets, with a focus on the key properties of the basis transformation and compression.
-In Section 5, we present the multiscale interpolation algorithm within the samplet basis and show an example of the resulting generalized Vandermonde matrix pattern.
+In Section 6, we present the multiscale interpolation algorithm in the samplet basis and provide error and cost estimates.
 Finally, in Section 7, we provide numerical tests in both two and three spatial dimensions.
 Concluding remarks are stated in Section 8.
+Throughout this article, to avoid the excessive use of generic constants, we use the notation $A\sim B$ to indicate that there exists constants $\underline{c},\overline{c}>0$ independent of $A,B$, and any quantities they may depend on, such that $\underline{c}A\leq B\leq\overline{c}A$.
+In addition $C>0$ will always denote a generic constant, whose value can change depending on the context.
 2. Classical multiscale method
-Let $\Omega\subseteq\mathbb{R}^{d}$ be a domain and $X=\{{\boldsymbol{x}}_{1},\ldots,{\boldsymbol{x}}_{N}\}\subset\Omega$ be a set of data sites with cardinality $N=\#X$.
-Moreover, let $f\in C(\Omega)$ be the target function generating the data $f_{1},\dots,f_{N}$.
-Associated to the set of sites $X$ there are two measures, the fill-distance of $X$ in $\Omega$ and the separation radius of $X$,
-$h_{X,\Omega}$ $\mathrel{\mathrel{\mathop{:}}=}\sup_{{\boldsymbol{x}}\in\Omega}\min_{{\boldsymbol{x}}\in X}\|{\boldsymbol{x}}-{\boldsymbol{x}}_{j}\|_{2}$
-$q_{X}$ $\mathrel{\mathrel{\mathop{:}}=}\frac{1}{2}\min_{j\neq k}\|{\boldsymbol{x}}_{j}-{\boldsymbol{x}}_{k}\|_{2}.$
-We call $X$ quasi-uniform if there is a constant $c_{\text{qu}}>0$ such that
-$q_{X}\leq h_{X,\Omega}\leq c_{\text{qu}}q_{X}.$
-Our goal is to interpolate or approximate the target function by means of radial basis functions (RBF) using only the given data $({\boldsymbol{x}}_{1},f_{1}),\ldots,({\boldsymbol{x}}_{N},f_{N})$.
+Let $\Omega\subseteq\mathbb{R}^{d}$ be a domain and $X=\{{\boldsymbol{x}}_{1},\ldots,{\boldsymbol{x}}_{N}\}\subseteq\Omega$ be a set of data sites with cardinality $N\mathrel{\mathrel{\mathop{:}}=}\#X$.
+Moreover, let $f\in C(\Omega)$ be the function generating the data values $f_{1},\dots,f_{N}$ corresponding to $\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{N}$.
+Associated to the set of data sites $X$, there are two measures, namely the fill-distance of $X$ in $\Omega$ and the separation radius of $X$, i.e.,
+$h_{X,\Omega}\mathrel{\mathrel{\mathop{:}}=}\sup_{{\boldsymbol{x}}\in\Omega}\min_{{\boldsymbol{x}}\in X}\|{\boldsymbol{x}}-{\boldsymbol{x}}_{j}\|_{2},\quad q_{X}\mathrel{\mathrel{\mathop{:}}=}\frac{1}{2}\min_{j\neq k}\|{\boldsymbol{x}}_{j}-{\boldsymbol{x}}_{k}\|_{2}.$
+We call $X$ quasi-uniform if there exists a constant $c_{\text{qu}}>0$ such that $q_{X}\leq h_{X,\Omega}\leq c_{\text{qu}}q_{X}.$
+Our goal is to interpolate or approximate the target function by means of RBFs only relying on the given data $({\boldsymbol{x}}_{1},f_{1}),\ldots,({\boldsymbol{x}}_{N},f_{N})$.
 A function $\Phi\colon\mathbb{R}^{d}\to\mathbb{R}$ is said to be radial, if and only if there exists a univariate function $\phi\colon[0,\infty)\to\mathbb{R}$ such that $\Phi({\boldsymbol{x}})=\phi(\|{\boldsymbol{x}}\|_{2})$.
-The radial function $\Phi$ is strictly positive definite, if and only if there holds for any choice of distinct points ${\boldsymbol{\xi}}_{1},\ldots,{\boldsymbol{\xi}}_{n}$, $n\in\mathbb{N}$, that
-${\boldsymbol{\alpha}}^{\intercal}[\Phi({\boldsymbol{\xi}}_{i}-{\boldsymbol{\xi}}_{j})]_{i,j=1}^{n}{\boldsymbol{\alpha}}>0\quad\text{for all }{\boldsymbol{\alpha}}\in\mathbb{R}^{n}\setminus\{{\boldsymbol{0}}\}.$
-In this case, the kernel
-$K({\boldsymbol{x}_{i}},{\boldsymbol{x}_{j}})\mathrel{\mathrel{\mathop{:}}=}\Phi({\boldsymbol{x}}_{i}-{\boldsymbol{x}}_{j})$
-is the reproducing kernel of a uniquely determined reproducing kernel Hilbert space, its native space, that we denote by $\big{(}\mathcal{N}_{\Phi},\langle\cdot,\cdot\rangle_{\mathcal{N}_{\Phi}}\big{)}$.
+The radial function $\Phi$ is strictly positive definite, if and only if there holds for any choice of distinct points ${\boldsymbol{\xi}}_{1},\ldots,{\boldsymbol{\xi}}_{n}$, $n\in\mathbb{N}$, that ${\boldsymbol{\alpha}}^{\intercal}[\Phi({\boldsymbol{\xi}}_{i}-{\boldsymbol{\xi}}_{j})]_{i,j}{\boldsymbol{\alpha}}>0\ \text{for all }{\boldsymbol{\alpha}}\in\mathbb{R}^{n}\setminus\{{\boldsymbol{0}}\}.$ In this case, the kernel $K({\boldsymbol{x}_{i}},{\boldsymbol{x}_{j}})\mathrel{\mathrel{\mathop{:}}=}\Phi({\boldsymbol{x}}_{i}-{\boldsymbol{x}}_{j})$ is the reproducing kernel of a uniquely determined reproducing kernel Hilbert space, its native space, that we denote by $\big{(}\mathcal{N}_{\Phi},\langle\cdot,\cdot\rangle_{\mathcal{N}_{\Phi}}\big{)}$.
 In this context, we will use the function $\Phi$ and its associated kernel $K$ synonymously.
-A popular class of strictly positive RBFs are the Matérn kernels or Sobolev splines $\Phi_{\nu}\colon\mathbb{R}^{d}\to\mathbb{R}$, dependent on a hyper-parameter $\nu>d/2$.
+A popular class of strictly positive RBFs are the Matérn kernels or Sobolev splines $\Phi_{\theta}\colon\mathbb{R}^{d}\to\mathbb{R}$, dependent on the hyper-parameter $\theta>d/2$, that steers the smoothness of the kernel.
 These kernels are defined via the radial function
-(1) $\phi_{\nu}(r)=\frac{2^{1-\nu}}{\Gamma(\nu)}r^{\nu-\frac{d}{2}}K_{\nu-\frac{d}{2}}(r),\quad r\geq 0,$
-where $\Gamma$ is the Riemann Gamma function and $K_{\beta}$ is the modified Bessel function of the second kind, see [12] for example.
-It is well-known that the Fourier transform of this radial function behaves like
-(2) $\widehat{\Phi}(\boldsymbol{\omega})=(1+\|\boldsymbol{\omega}\|_{2}^{2})^{-\nu-\frac{d}{2}},\quad\boldsymbol{\omega}\in\mathbb{R}^{d},$
-and that, in turn, the native space of the corresponding Matérn kernel is norm-equivalent to the classical Sobolev space $H^{\nu+\frac{d}{2}}(\mathbb{R}^{d})$, see [20, Theorem 6.13] for example.
-In this instationary setting it is known that the interpolant
-$s\in\mathcal{S}\mathrel{\mathrel{\mathop{:}}=}\operatorname{span}\{\Phi(\cdot-{\boldsymbol{x}}):{\boldsymbol{x}}\in X\}\subset\mathcal{N}_{\Phi}$
-converges towards the target function $f$ as $h_{X,\Omega}\to 0$.
-However, this improved approximation comes with higher and higher numerical cost for the iterative solution due to the ill-conditioning.
+(1) $\phi_{\theta}(r)=\frac{2^{1-\theta+d/2}}{\Gamma(\theta-d/2)}r^{\theta-d/2}K_{\theta-d/2}(r),\quad r\geq 0,$
+where $\Gamma$ is the Riemann gamma function and $K_{\nu}$ is the modified Bessel function of the second kind of order $\nu$, see [19] for example.
+It is well-known that the Fourier transform of this radial function satisfies
+(2) $\mathcal{F}{\Phi}(\boldsymbol{\omega})=(1+\|\boldsymbol{\omega}\|_{2}^{2})^{-\theta},\quad\boldsymbol{\omega}\in\mathbb{R}^{d},$
+and that, in turn, the native space of the corresponding Matérn kernel is norm-equivalent to the Sobolev space $H^{\theta}(\mathbb{R}^{d})$, see [28, Corollary 10.13] for example.
+In this instationary setting, meaning that the length-scale of the kernel is not coupled to the fill-distance, it is known that the interpolant $s\in\mathcal{S}\mathrel{\mathrel{\mathop{:}}=}\operatorname{span}\{\Phi(\cdot-{\boldsymbol{x}}):{\boldsymbol{x}}\in X\}\subset\mathcal{N}_{\Phi}$ converges towards the function $f$ as $h_{X,\Omega}\to 0$.
+However, this improved approximation comes with higher and higher numerical cost for the iterative solution due to the ill-conditioning of the associated generalized Vandermonde matrix.
 To alleviate this cost, we can rescale the kernel, i.e., we introduce the parameter $\delta>0$ and define
 (3) $\Phi_{\delta}\mathrel{\mathrel{\mathop{:}}=}\delta^{-d}\Phi\bigg{(}\frac{\cdot}{\delta}\bigg{)}.$
-In the stationary setting we couple $\delta$ proportional to $h_{X,\Omega}$, i.e., there is a coupling parameter $\eta>0$ such that $\delta=\eta h_{X,\Omega}$.
-In this case we can control the numerical costs to compute the interpolant but cannot expect convergence.
-This is known as the trade-off principle.
+In this stationary setting, we couple $\delta$ proportional to $h_{X,\Omega}$, i.e., there is a coupling parameter $\eta>0$ such that
+(4) $\delta=\eta h_{X,\Omega}.$
+In this case, we can control the numerical costs to compute the interpolant but cannot expect convergence. This is known as the trade-off principle.
 The idea of the multiscale method is now to introduce a hierarchy
 (5) $X_{1},X_{2},\ldots,X_{L}=X,$
-where we refer to the subscript of $X_{\ell}$ as level with cardinalities $N_{\ell}$ and fill-distances $h_{\ell}\mathrel{\mathrel{\mathop{:}}=}h_{X_{\ell},\Omega}$.
+where we refer to the subscript of $X_{\ell}$ as level with cardinalities $N_{\ell}\mathrel{\mathrel{\mathop{:}}=}\#X_{\ell}$ and fill-distances $h_{\ell}\mathrel{\mathrel{\mathop{:}}=}h_{X_{\ell},\Omega}$.
 Although in many applications the sets will be nested, for the multiscale method itself, this is not necessary.
@@
 We quantify this relation in Theorem 2.1.
-Now, we use the stationary setting on each level $\ell$.
-We have level-dependent $\delta_{\ell}>0$ and level-dependently rescaled kernel $K_{\ell}$ given by
-$K_{\ell}(\boldsymbol{x},\boldsymbol{y})\mathrel{\mathrel{\mathop{:}}=}\Phi_{\ell}(\boldsymbol{x}-\boldsymbol{y})=\delta_{\ell}^{-d}\Phi\left(\frac{\boldsymbol{x}-\boldsymbol{y}}{\delta_{\ell}}\right),\quad\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d}.$
+In what follows, we consider the stationary setting at each level $\ell$.
+This yields level-dependent parameters $\delta_{\ell}>0$ and level-dependently rescaled kernels $K_{\ell}$ given by
+(6) $K_{\ell}(\boldsymbol{x},\boldsymbol{y})\mathrel{\mathrel{\mathop{:}}=}\Phi_{\ell}(\boldsymbol{x}-\boldsymbol{y})=\delta_{\ell}^{-d}\Phi\left(\frac{\boldsymbol{x}-\boldsymbol{y}}{\delta_{\ell}}\right),\quad\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d},\ \ell=1,\ldots,L.$
 Each set $X_{\ell}$ in the multiscale hierarchy (5), together with the respective kernel, gives rise to a subspace
 (7) $\mathcal{S}_{\ell}\mathrel{\mathrel{\mathop{:}}=}\operatorname{span}\{\Phi_{\ell}(\cdot-{\boldsymbol{x}}):{\boldsymbol{x}}\in X_{\ell}\}\subset\mathcal{N}_{\Phi_{\ell}},$
-also called local approximation space.
-With the multiscale sequence of points and the associated kernels, we can decompose the original approximation problem into $L$ steps using a telescoping sum
+called the local approximation space.
+Now, given the multiscale sequence $X_{1},\ldots,X_{L}$, we seek an approximation in terms of successive residual corrections. To this end, we introduce functions $w_{\ell}$ of the form
+(8) $w_{\ell}=\sum_{i=1}^{N_{\ell}}c_{\ell,i}\Phi_{\ell}(\cdot-{\boldsymbol{x}}_{i})\in\mathcal{S}_{\ell}.$
+The coefficents $c_{\ell,i}$ are determined from decomposing the approximation problem by a telescoping sum according to
 $\left.s_{1}\right|_{X_{1}}\mathrel{=\mathrel{\mathop{:}}}\left.w_{1}\right|_{X_{1}}$ $=\left.f\right|_{X_{1}},$
@@
 $w_{L}|_{X_{L}}$ $=\bigg{(}f-\sum_{\ell=1}^{L-1}w_{\ell}\bigg{)}\bigg{|}_{X_{L}}.$
-The intuition is that at the $\ell$-th step we compute the approximation to the error function or residual of level $\ell-1$, i.e.,
-(6) $f-(w_{1}+\cdots+w_{l-1})$
-in the space $\mathcal{S}_{\ell}$, given as
-(7) $w_{\ell}({\boldsymbol{x}})=\sum_{i=1}^{N_{\ell}}c_{\ell,i}\Phi_{\ell}({\boldsymbol{x}}-{\boldsymbol{x}}_{i}),\quad\boldsymbol{x}\in\Omega,$
-with a coefficient vector ${\boldsymbol{c}}_{\ell}\in\mathbb{R}^{N_{\ell}}$.
-From this, we obtain the multiscale decomposition
-$s_{L}\mathrel{\mathrel{\mathop{:}}=}(w_{1}+\cdots+w_{L})|_{X}=f|_{X},$
-which means that $s_{L}\in\cup_{\ell=1}^{L}\mathcal{S}_{\ell}$ matches $f$ at each of the data sites in $X$.
-In practice, the multiscale approximation process begins with a coarse set of points, where the approximation problem (7) is solved.
+This means that, at level $\ell$, we compute the interpolant to the residual at level $\ell-1$, i.e.,
+(9) $f-(w_{1}+\cdots+w_{\ell-1})$
+in the space $\mathcal{S}_{\ell}$.
+From this, we obtain the multiscale decomposition $s_{L}\mathrel{\mathrel{\mathop{:}}=}(w_{1}+\cdots+w_{L})\in\cup_{\ell=1}^{L}\mathcal{S}_{\ell}$ which matches $f$ at each of the data sites in $X$, i.e., $s_{L}|_{X}=f|_{X}$.
+In practice, the multiscale approximation process begins with a small set of points, where the approximation problem (8) is solved.
 Next, the error function (9) is computed, and finally, the latter is approximated at a finer scale.
-Interpolating the residual at level $\ell$ can be written in the following form:
-$\boldsymbol{K}_{\ell,\ell}\boldsymbol{c}_{\ell}=\boldsymbol{f}_{\ell}-\sum_{\ell^{\prime}=1}^{\ell-1}\boldsymbol{K}_{\ell,\ell^{\prime}}\boldsymbol{c}_{\ell^{\prime}},$
-where
-$\boldsymbol{K}_{\ell,\ell^{\prime}}\mathrel{\mathrel{\mathop{:}}=}\big{[}\Phi_{\ell^{\prime}}\big{(}{\boldsymbol{x}}_{i}^{(\ell)}-\boldsymbol{x}_{j}^{(\ell^{\prime})}\big{)}\big{]}_{i,j},\quad\ell,\ell^{\prime}=1,\ldots,L.$
-Therefore, the algorithm to solve the multiscale interpolation consists of solving the following triangular system, whose entries are matrices of different sizes:
-(8) $\left[\begin{array}[]{ccccc}\boldsymbol{K}_{1,1}&&&&\\ \boldsymbol{K}_{2,1}&\boldsymbol{K}_{2,2}&&&\\ \boldsymbol{K}_{3,1}&\boldsymbol{K}_{3,2}&\boldsymbol{K}_{3,3}&&\\ \vdots&\vdots&\cdots&\ddots&\\ \boldsymbol{K}_{L,1}&\boldsymbol{K}_{L,2}&\cdots&\boldsymbol{K}_{L,L-1}&\boldsymbol{K}_{L,L}\end{array}\right]\left[\begin{array}[]{c}\boldsymbol{c}_{1}\\ \boldsymbol{c}_{2}\\ \boldsymbol{c}_{3}\\ \vdots\\ \boldsymbol{c}_{L}\end{array}\right]=\left[\begin{array}[]{c}\boldsymbol{f}_{1}\\ \boldsymbol{f}_{2}\\ \boldsymbol{f}_{3}\\ \vdots\\ \boldsymbol{f}_{L}\end{array}\right]$
-The resulting matrix is in general a densely-populated populated lower triangular matrix as a result of the choice of globally supported RBFs for the approximation at each level.
-The samplet compression plays a crucial role here, rendering the matrix quasi-sparse.
-This technique is introduced in Section 4, and the linear system is reformulated in the samplet basis in Section 5.
+The coefficients of the residual correction at level $\ell$ can be computed by solving
+$\boldsymbol{K}_{\ell,\ell}\boldsymbol{c}_{\ell}=\boldsymbol{f}_{\ell}-\sum_{\ell^{\prime}=1}^{\ell-1}\boldsymbol{K}_{\ell,\ell^{\prime}}\boldsymbol{c}_{\ell^{\prime}}$
+with the generalized Vandermonde matrices $\boldsymbol{K}_{\ell,\ell^{\prime}}\mathrel{\mathrel{\mathop{:}}=}\big{[}\Phi_{\ell^{\prime}}\big{(}{\boldsymbol{x}}_{i}^{(\ell)}-\boldsymbol{x}_{j}^{(\ell^{\prime})}\big{)}\big{]}_{i,j}\in\mathbb{R}^{N_{\ell}\times N_{\ell^{\prime}}}$ and the vectors $\boldsymbol{f}_{\ell}\mathrel{\mathrel{\mathop{:}}=}\boldsymbol{f}|_{X_{\ell}}\in\mathbb{R}^{N_{\ell}},\ \boldsymbol{c}_{\ell}\mathrel{\mathrel{\mathop{:}}=}[{c}_{\ell,i}]_{i}\in\mathbb{R}^{N_{\ell}}.$ Consequently, the coefficient vectors ${\boldsymbol{c}}_{1},\ldots,{\boldsymbol{c}}_{L}$ are obtained by solving the block lower triangular system
+(10) $\mathcal{K}{\boldsymbol{c}}\mathrel{\mathrel{\mathop{:}}=}\left[\begin{array}[]{cccc}\boldsymbol{K}_{1,1}&&&\\ \boldsymbol{K}_{2,1}&\boldsymbol{K}_{2,2}&&\\ \vdots&\vdots&\ddots&\\ \boldsymbol{K}_{L,1}&\boldsymbol{K}_{L,2}&\cdots&\boldsymbol{K}_{L,L}\end{array}\right]\left[\begin{array}[]{c}\boldsymbol{c}_{1}\\ \boldsymbol{c}_{2}\\ \vdots\\ \boldsymbol{c}_{L}\end{array}\right]=\left[\begin{array}[]{c}\boldsymbol{f}_{1}\\ \boldsymbol{f}_{2}\\ \vdots\\ \boldsymbol{f}_{L}\end{array}\right]\mathrel{=\mathrel{\mathop{:}}}{\boldsymbol{f}}.$
+The resulting block matrix $\mathcal{K}$ is in general a densely-populated lower triangular matrix as a result of the choice of globally supported RBFs for the approximation at each level.
 There holds the following convergence result, see, e.g., [29, 30].
@@
 for $\ell=1,2,\dots$ with fixed $\mu\in(0,1)$, $c\in(0,1]$ and $h_{1}$ sufficiently small.
-Let $\Phi\colon\mathbb{R}^{d}\to\mathbb{R}$ be a reproducing kernel for $H^{s}(\mathbb{R}^{d})$, i.e., its Fourier transform satisfies
-$c_{1}(1+\|{\boldsymbol{\omega}}\|_{2}^{2})^{-s}\leq\widehat{\Phi}({\boldsymbol{\omega}})\leq c_{2}(1+\|{\boldsymbol{\omega}}\|_{2}^{2})^{-s},\quad{\boldsymbol{\omega}}\in\mathbb{R}^{d},$
-and let $\Phi_{\ell}$ be the rescaled kernel as in (3) with scale factor $\delta_{\ell}=\eta h_{\ell}$, $\eta>0$. Let $f\in H^{s}(\Omega)$. Then there exist constants $C$ and $C_{1}>0$ such that
-$\|f-s_{L}\|_{L_{2}(\Omega)}\leq C(C_{1}\mu^{s}+C_{1}\eta^{-s})^{L}\|f\|_{H^{s}(\Omega)}.$
+Further, let $\Phi\colon\mathbb{R}^{d}\to\mathbb{R}$ be a reproducing kernel for $H^{\theta}(\mathbb{R}^{d})$, i.e., its Fourier transform satisfies $\mathcal{F}{\Phi}({\boldsymbol{\omega}})\sim(1+\|{\boldsymbol{\omega}}\|_{2}^{2})^{-\theta},\ {\boldsymbol{\omega}}\in\mathbb{R}^{d}$ and let $\Phi_{\ell}$ be the rescaled kernel as in (3) with scale factor $\delta_{\ell}=\eta h_{\ell}$, $\eta>1$.
+Let $f\in H^{\theta}(\Omega)$.
+Then there exist constants $C$ and $C_{1}>0$ such that
+$\|f-s_{L}\|_{L_{2}(\Omega)}\leq C(C_{1}\mu^{\theta}+C_{1}\eta^{-\theta})^{L}\|f\|_{H^{\theta}(\Omega)}.$
 3. Condition number bound for generalized Vandermonde matrices
-We now show that the condition number $\operatorname{cond}_{2}$ of the generalized Vandermonde matrix $\boldsymbol{K}_{\ell,\ell}$ of the scaled kernel can be bounded by a constant, independently of the level $\ell$.
+We now show that the condition number of the generalized Vandermonde matrices $\boldsymbol{K}_{\ell,\ell}$, $\ell=1,\ldots,L$, of the scaled kernel can be bounded by a constant, independently of the level $\ell$.
 A similar result in the context of the multiscale method for compactly supported RBFs has been derived in [26] using techniques of [22, 24].
@@
 Furthermore, for a lighter notation, we will drop the subscripts $\ell$ in this section.
-It goes without saying that, such a bound can not be established for general sets of sites $X$.
+It goes without saying that, such a bound can not be established for general sets of data sites $X$.
 Hence, for this section, we assume that $X$ is quasi-uniform and the parameters of the multilevel method are chosen as in Theorem 2.1.
-We remark that the algebraic decay of the Fourier transform of the Matérn kernel (2) immediately yields for the scaled RBF $\Phi_{\delta}$ that
-(10) $\widehat{\Phi_{\delta}}(\boldsymbol{\omega})=(1+\delta^{2}\|\boldsymbol{\omega}\|_{2}^{2})^{-\nu-\frac{d}{2}},\quad\boldsymbol{\omega}\in\mathbb{R}^{d}.$
-We will bound the smallest and largest eigenvalue of the matrix ${\boldsymbol{K}}$ separately. We follow the ideas of [18, 15, 16].
+We remark that the algebraic decay of the Fourier transform of the Matérn kernel (2) immediately yields for the scaled RBF $\Phi_{\delta}$ that $\mathcal{F}{\Phi_{\delta}}(\boldsymbol{\omega})=(1+\delta^{2}\|\boldsymbol{\omega}\|_{2}^{2})^{-\theta},\ \boldsymbol{\omega}\in\mathbb{R}^{d}.$ We will bound the smallest and largest eigenvalue of the matrix ${\boldsymbol{K}}$ separately.
+We follow the ideas of [26, 22, 24].
 Theorem 3.1.
-Let $\Omega\subseteq\mathbb{R}^{d}$ be a bounded domain and $X\subseteq\Omega$ be a quasi-uniform set of sites with fill-distance $h_{X,\Omega}$ and separation radius $q_{X}$.
-Let $\Phi$ be the Matérn kernel and Let $\Phi_{\delta}=\delta^{-d}\Phi(\cdot/\delta)$ be the rescaled kernel where $\delta=\eta h_{X,\Omega}$ with overlap parameter $\eta>1$.
-Then there exists a constant $C=C(\Phi,d,s)$ such that the bound for the smallest eigenvalue $\lambda_{min}$ of $\boldsymbol{K}=[\Phi_{\delta}(\boldsymbol{x}_{i}-\boldsymbol{x}_{j})]_{i,j}$
-$\lambda_{min}({\boldsymbol{K}})\geq C(\eta c_{\textnormal{qu}})^{d-2s}\delta^{-d}$
-holds. The largest eigenvalue of ${\boldsymbol{K}}$ can be bounded from above by
-$\lambda_{max}({\boldsymbol{K}})\leq\widetilde{C}\delta^{-d},$
-with a constant $\widetilde{C}=C(\Phi,d,\nu,\eta,c_{\textnormal{qu}})$.
+Let $\Omega\subseteq\mathbb{R}^{d}$ be a bounded domain and $X\subseteq\Omega$ be a quasi-uniform set of data sites with fill-distance $h_{X,\Omega}$ and separation radius $q_{X}$.
+Let $\Phi$ be the Matérn kernel with smoothness parameter $\theta$.
+Let $\Phi_{\delta}=\delta^{-d}\Phi(\cdot/\delta)$ be the rescaled kernel where $\delta=\eta h_{X,\Omega}$ with overlap parameter $\eta>1$.
+Then there exists a constant $C={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}C(\Phi,d)}$ such that the the smallest eigenvalue $\lambda_{\min}$ of $\boldsymbol{K}=[\Phi_{\delta}(\boldsymbol{x}_{i}-\boldsymbol{x}_{j})]_{i,j}$ satisfies $\lambda_{\min}({\boldsymbol{K}})\geq C(\eta c_{\textnormal{qu}})^{d-2\theta}\delta^{-d},$ while the largest eigenvalue of ${\boldsymbol{K}}$ can be bounded from above by $\lambda_{\max}({\boldsymbol{K}})\leq\widetilde{C}\delta^{-d},$ with a constant $\widetilde{C}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}C(\Phi,d,\eta,c_{\textnormal{qu}})}$.
 Proof.
@@
 $\widetilde{\boldsymbol{K}}=\delta^{d}{\boldsymbol{K}}=\bigg{[}\Phi\bigg{(}\frac{{\boldsymbol{x}}_{i}-{\boldsymbol{x}}_{j}}{\delta}\bigg{)}\bigg{]}_{i,j}.$
-The lower bound follows immediately from [18, Proof of Proposition 3.4].
-We only need to prove the upper bound for $\lambda_{max}({\boldsymbol{K}})$.
-We use Gershgorin’s circle theorem, see, e.g.
-[23, Theorem 5.1] for $\widetilde{\boldsymbol{K}}$.
-Since the diagonal elements in this matrix are $\Phi(\boldsymbol{0})$, independent of the line $j$, we can, without loss of generality, set $j=1$ and $\boldsymbol{x}_{1}=\boldsymbol{0}$.
+The lower bound immediately follows from [26, Proof of Proposition 3.4].
+Hence, we only prove the upper bound for $\lambda_{\max}({\boldsymbol{K}})$.
+To this end, we apply Gershgorin’s circle theorem, see, e.g., [31, Theorem 5.1] to $\widetilde{\boldsymbol{K}}$.
+Since the diagonal elements of this matrix are $\Phi(\boldsymbol{0})$, independently of the row $i$, we can, without loss of generality, set $i=1$ and $\boldsymbol{x}_{1}=\boldsymbol{0}$.
 We have
@@
 #$\#\{\boldsymbol{x}_{k}\in E_{n}\}\leq 3^{d}n^{d-1},$
-see, e.g., [20, Proof of Theorem 12.3]. Inserting this into (11) yields
-$|\lambda_{max}(\widetilde{\boldsymbol{K}})-\Phi(\boldsymbol{0})|$ $\leq\sum_{k=2}^{N}\bigg{|}\Phi\bigg{(}\frac{\boldsymbol{x}_{k}}{\delta}\bigg{)}\bigg{|}$
-$\leq\sum_{n=1}^{\infty}\sum_{\boldsymbol{x}_{k}\in E_{n}}\bigg{|}\Phi\bigg{(}\frac{\boldsymbol{x}_{k}}{\delta}\bigg{)}\bigg{|}$
-(12) $\leq\sum_{n=1}^{\infty}\#\{\boldsymbol{x}_{k}\in E_{n}\}\sup_{\boldsymbol{x}\in E_{n}}\bigg{|}\Phi\bigg{(}\frac{\boldsymbol{x}}{\delta}\bigg{)}\bigg{|}.$
+see, e.g., [28, Proof of Theorem 12.3]. Inserting this into (12) yields
+(13) $|\lambda_{\max}(\widetilde{\boldsymbol{K}})-\Phi(\boldsymbol{0})|$ $\leq\sum_{k=2}^{N}\bigg{|}\Phi\bigg{(}\frac{\boldsymbol{x}_{k}}{\delta}\bigg{)}\bigg{|}\leq\sum_{n=1}^{\infty}\sum_{\boldsymbol{x}_{k}\in E_{n}}\bigg{|}\Phi\bigg{(}\frac{\boldsymbol{x}_{k}}{\delta}\bigg{)}\bigg{|}$
+$\leq\sum_{n=1}^{\infty}\#\{\boldsymbol{x}_{k}\in E_{n}\}\sup_{\boldsymbol{x}\in E_{n}}\bigg{|}\Phi\bigg{(}\frac{\boldsymbol{x}}{\delta}\bigg{)}\bigg{|}.$
 Next, we bound $\sup_{\boldsymbol{x}\in E_{n}}|\Phi({\boldsymbol{x}}/{\delta})|$.
 We use the specific form of $\Phi$, cp.
-(1), and the fact that the function $r\mapsto r^{\beta}K_{\beta}(r)$ is non-increasing on $(0,\infty)$ for any $\beta\in\mathbb{R}$, see [20, Corollary 5.12].
+(1), and the fact that the function $r\mapsto r^{\beta}K_{\beta}(r)$ is non-increasing on $(0,\infty)$ for any $\beta\in\mathbb{R}$, see [28, Corollary 5.12].
 This yields
-$\frac{\Gamma(\nu)}{2^{1-\nu}}\sup_{\boldsymbol{x}\in E_{n}}\bigg{|}\Phi\bigg{(}\frac{\boldsymbol{x}}{\delta}\bigg{)}\bigg{|}$ $=\sup_{\boldsymbol{x}\in E_{n}}\bigg{(}\frac{\|\boldsymbol{x}\|_{2}}{\delta}\bigg{)}^{\nu-\frac{d}{2}}K_{\nu-\frac{d}{2}}\bigg{(}\frac{\|\boldsymbol{x}\|_{2}}{\delta}\bigg{)}$
-$\leq\bigg{(}\frac{nq_{X}}{\delta}\bigg{)}^{\nu-\frac{d}{2}}K_{\nu-\frac{d}{2}}\bigg{(}\frac{nq_{X}}{\delta}\bigg{)}$
-$\leq\bigg{(}\frac{nq_{X}}{\delta}\bigg{)}^{\nu-\frac{d}{2}}\sqrt{2\pi}\bigg{(}\frac{\delta}{nq_{X}}\bigg{)}^{\frac{1}{2}}\exp\bigg{(}-\frac{nq_{X}}{\delta}\bigg{)}\exp\Bigg{(}\bigg{(}\nu-\frac{d}{2}\bigg{)}^{2}\frac{2\delta}{2nq_{X}}\Bigg{)},$
-where we used the asymptotic behavior [20, Lemma 5.13] of the modified Bessel function to arrive at the last inequality.
+$\frac{\Gamma\left(\theta-\frac{d}{2}\right)}{2^{1-\theta+\frac{d}{2}}}\sup_{\boldsymbol{x}\in E_{n}}\bigg{|}\Phi\bigg{(}\frac{\boldsymbol{x}}{\delta}\bigg{)}\bigg{|}=\sup_{\boldsymbol{x}\in E_{n}}\bigg{(}\frac{\|\boldsymbol{x}\|_{2}}{\delta}\bigg{)}^{\theta-\frac{d}{2}}K_{\theta-\frac{d}{2}}\bigg{(}\frac{\|\boldsymbol{x}\|_{2}}{\delta}\bigg{)}$
+$\leq\bigg{(}\frac{nq_{X}}{\delta}\bigg{)}^{\theta-\frac{d}{2}}K_{\theta-\frac{d}{2}}\bigg{(}\frac{nq_{X}}{\delta}\bigg{)}$
+$\leq\bigg{(}\frac{nq_{X}}{\delta}\bigg{)}^{\theta-\frac{d}{2}}\sqrt{2\pi}\bigg{(}\frac{\delta}{nq_{X}}\bigg{)}^{\frac{1}{2}}\exp\bigg{(}-\frac{nq_{X}}{\delta}\bigg{)}\exp\Bigg{(}\bigg{(}\theta-\frac{d}{2}\bigg{)}^{2}\frac{2\delta}{2nq_{X}}\Bigg{)},$
+where we used the asymptotic behavior [28, Lemma 5.13] of the modified Bessel function to arrive at the last inequality.
 Inserting this into (13), using the quasi-uniformity of $X$ and the coupling of $\delta$ to $h_{X,\Omega}$, as in Theorem 2.1, yields
-$|\lambda_{max}(\widetilde{\boldsymbol{K}})-\Phi(\boldsymbol{0})|$ $\leq\sum_{n=1}^{\infty}\#\{\boldsymbol{x}_{k}\in E_{n}\}\sup_{\boldsymbol{x}\in E_{n}}\bigg{|}\Phi\bigg{(}\frac{\boldsymbol{x}}{\delta}\bigg{)}\bigg{|}$
-$\leq\sum_{n=1}^{\infty}3^{d}n^{d-1}\sqrt{2\pi}\eta^{\frac{d}{2}-\nu}(\eta c_{\text{qu}})^{\frac{1}{2}}n^{\nu-\frac{d}{2}-\frac{1}{2}}\exp\bigg{(}-\frac{n}{\eta}+\frac{(\nu-\frac{d}{2})^{2}\eta c_{q}u}{2n}\bigg{)}$
-$\leq\sqrt{2\pi}3^{d}\eta^{\frac{d}{2}-\nu}(\eta c_{\text{qu}})^{\frac{1}{2}}\sum_{n=1}^{\infty}n^{\nu-\frac{d}{2}-\frac{3}{2}}\exp\bigg{(}-\frac{n}{\eta}+\frac{(\nu-\frac{d}{2})^{2}\eta c_{q}u}{2n}\bigg{)}.$
-Since the series on the right hand side of the estimate converges, this yields the claimed bound for $\lambda_{max}$. ∎
-The derived bounds on $\lambda_{min}$ and $\lambda_{max}$ lead to the boundedness of the condition number of the generalized Vandermonde matrix for every level, uniformly in the number of data sites and the scaling parameter.
+$|\lambda_{\max}(\widetilde{\boldsymbol{K}})-\Phi(\boldsymbol{0})|$ $\leq\sum_{n=1}^{\infty}\#\{\boldsymbol{x}_{k}\in E_{n}\}\sup_{\boldsymbol{x}\in E_{n}}\bigg{|}\Phi\bigg{(}\frac{\boldsymbol{x}}{\delta}\bigg{)}\bigg{|}$
+$\leq\sum_{n=1}^{\infty}3^{d}n^{d-1}\sqrt{2\pi}\eta^{\frac{d}{2}-\theta}(\eta c_{\text{qu}})^{\frac{1}{2}}n^{\theta-\frac{d}{2}-\frac{1}{2}}\exp\bigg{(}-\frac{n}{\eta}+\frac{\left(\theta-\frac{d}{2}\right)^{2}\eta c_{\text{qu}}}{n}\bigg{)}$
+$\leq\sqrt{2\pi}3^{d}\eta^{\frac{d}{2}-\theta}(\eta c_{\text{qu}})^{\frac{1}{2}}\sum_{n=1}^{\infty}n^{\theta-\frac{d}{2}-\frac{3}{2}}\exp\bigg{(}-\frac{n}{\eta}+\frac{\left(\theta-\frac{d}{2}\right)^{2}\eta c_{\text{qu}}}{n}\bigg{)}.$
+Since the series on the right hand side of the estimate converges, this yields the claimed bound for $\lambda_{\max}$. ∎
+The derived bounds on $\lambda_{\min}$ and $\lambda_{\max}$ lead to the boundedness of the condition number of the generalized Vandermonde matrices ${\boldsymbol{K}}_{\ell,\ell}$ for $\ell=1,\ldots,L$, uniformly in the number of data sites and the scaling parameter.
+In addition, by a suitable diagonal scaling, also the lower triangular system in (10) becomes well conditioned.
+We have the following
 Corollary 3.2.
 For every $\ell\in\mathbb{N}$ the condition number $\operatorname{cond}_{2}(\boldsymbol{K}_{\ell,\ell})$ is bounded independently of $N_{\ell}$ and $\delta_{\ell}$.
+Furthermore, introducing the block diagonal scaling $\mathcal{D}\mathrel{\mathrel{\mathop{:}}=}\operatorname{diag}(\delta_{1}^{d}{\boldsymbol{I}},\ldots,\delta_{L}^{d}{\boldsymbol{I}}),$ the diagonally scaled matrix $\mathcal{D}^{1/2}\mathcal{K}\mathcal{D}^{1/2}$, cp.
+(10) is well conditioned, i.e., the corresponding condition number satisfies $\operatorname{cond}_{2}(\mathcal{D}^{1/2}\mathcal{K}\mathcal{D}^{1/2})\sim\operatorname{cond}_{F}(\mathcal{D}^{1/2}\mathcal{K}\mathcal{D}^{1/2})\sim 1$.
+4. Numerical approximation of the multiscale system
+In this paragraph, we discuss the approximation error issuing from replacing the matrix $\mathcal{K}$ of the multiscale system in (10) by an approximation $\widehat{\mathcal{K}}$.
+As has been observed in Corollary 3.2, the multiscale system becomes well-conditioned by suitable diagonal scaling.
+For our analysis, we therefore consider the preconditioned matrix $\mathcal{K}_{\mathrm{pc}}\mathrel{\mathrel{\mathop{:}}=}\mathcal{D}^{1/2}\mathcal{K}\mathcal{D}^{1/2}$ and its numerical approximation $\widehat{\mathcal{K}}_{\mathrm{pc}}\mathrel{\mathrel{\mathop{:}}=}\mathcal{D}^{1/2}\widehat{\mathcal{K}}\mathcal{D}^{1/2}$.
+We have the following standard perturbation result, see, e.g., [12].
+Lemma 4.1.
+Let $\widehat{\mathcal{K}}$ be sufficiently close to $\mathcal{K}$ and let $\mathcal{K}_{\mathrm{pc}}{\boldsymbol{d}}=\mathcal{D}^{1/2}{\boldsymbol{f}}$ and $\widehat{\mathcal{K}}_{\mathrm{pc}}\widehat{\boldsymbol{d}}=\mathcal{D}^{1/2}{\boldsymbol{f}}$.
+Then there exists a constant $C>0$ such that the relative consistency error satisfies
+$\frac{\|{\boldsymbol{d}}-\widehat{\boldsymbol{d}}\|_{2}}{\|{\boldsymbol{d}}\|_{2}}\leq C\frac{\|\mathcal{K}_{\mathrm{pc}}-\widehat{\mathcal{K}}_{\mathrm{pc}}\|_{F}}{\|\mathcal{K}_{\mathrm{pc}}\|_{F}}.$
+Proof.
+The proof follows similarly to the proof of [12, Theorem 7.2]. There holds
+$\mathcal{K}_{\mathrm{pc}}(\widehat{\boldsymbol{d}}-{\boldsymbol{d}})=-(\widehat{\mathcal{K}}_{\mathrm{pc}}-\mathcal{K}_{\mathrm{pc}}){\boldsymbol{d}}+(\widehat{\mathcal{K}}_{\mathrm{pc}}-\mathcal{K}_{\mathrm{pc}})({\boldsymbol{d}}-\widehat{\boldsymbol{d}})$
+Hence, by multiplying by $\mathcal{K}_{\mathrm{pc}}^{-1}$, taking norms and exploiting the consistency of the Frobenius and the Euclidean norm, we obtain
+$\|{\boldsymbol{d}}-\widehat{\boldsymbol{d}}\|_{2}\leq\|\mathcal{K}_{\mathrm{pc}}^{-1}\|_{F}\|\widehat{\mathcal{K}}_{\mathrm{pc}}-\mathcal{K}_{\mathrm{pc}}\|_{F}\|{\boldsymbol{d}}\|_{2}+\|\mathcal{K}_{\mathrm{pc}}^{-1}\|_{F}\|\widehat{\mathcal{K}}_{\mathrm{pc}}-\mathcal{K}_{\mathrm{pc}}\|_{F}\|{\boldsymbol{d}}-\widehat{\boldsymbol{d}}\|_{2}.$
+Factoring out common terms and simplifying the expression then results in
+$\|{\boldsymbol{d}}-\widehat{\boldsymbol{d}}\|_{2}\leq\frac{\|\mathcal{K}_{\mathrm{pc}}^{-1}\|_{F}\|\widehat{\mathcal{K}}_{\mathrm{pc}}-\mathcal{K}_{\mathrm{pc}}\|_{F}}{1-\|\mathcal{K}_{\mathrm{pc}}^{-1}\|_{F}\|\widehat{\mathcal{K}}_{\mathrm{pc}}-\mathcal{K}_{\mathrm{pc}}\|_{F}}\|{\boldsymbol{d}}\|_{2}.$
+Dividing by $\|{\boldsymbol{d}}\|_{2}$ and expanding the respective terms in the numerator and denominator by $\|\mathcal{K}_{\mathrm{pc}}\|_{F}$ finally gives
+$\frac{\|{\boldsymbol{d}}-\widehat{\boldsymbol{d}}\|_{2}}{\|{\boldsymbol{d}}\|_{2}}\leq\frac{\operatorname{cond}_{F}(\mathcal{K}_{\mathrm{pc}})}{1-\operatorname{cond}_{F}(\mathcal{K}_{\mathrm{pc}})\|\mathcal{K}_{\mathrm{pc}}-\widehat{\mathcal{K}}_{\mathrm{pc}}\|_{F}/\|\mathcal{K}_{\mathrm{pc}}\|_{F}}\frac{\|\mathcal{K}_{\mathrm{pc}}-\widehat{\mathcal{K}}_{\mathrm{pc}}\|_{F}}{\|\mathcal{K}_{\mathrm{pc}}\|_{F}}.$
+For a sufficiently small consistency error, we arrive at the claimed bound using Corollary 3.2. ∎
+To connect the previous error estimate to the consistency error of the multiscale interpolant, we require the following technical lemma.
+Lemma 4.2.
+Consider any function $v_{\ell}=\sum_{i=1}^{N_{\ell}}a_{i}\Phi_{\ell}(\cdot-{\boldsymbol{x}}_{i}^{(\ell)})\in\mathcal{S}_{\ell}$. Then, there holds
+$\|v_{\ell}\|_{\mathcal{N}_{\Phi_{\ell}}}\sim\delta_{\ell}^{-d/2}\|{\boldsymbol{a}}\|_{2}\quad\text{as well as}\quad\|v_{\ell}|X_{\ell}\|_{\ell^{2}}\sim\delta_{\ell}^{-d}\|{\boldsymbol{a}}\|_{2},$
+where ${\boldsymbol{a}}\mathrel{\mathrel{\mathop{:}}=}[a_{1},\ldots,a_{N_{\ell}}]^{\intercal}$ is the coefficient vector of $v_{\ell}$.
+Proof.
+By the reproducing property and due to $\|{\boldsymbol{K}}_{\ell,\ell}\|_{2}\sim\|{\boldsymbol{K}}_{\ell,\ell}^{-1}\|_{2}\sim\delta^{-d}_{\ell}$, we obtain
+$\|v_{\ell}\|_{\mathcal{N}_{\Phi_{\ell}}}^{2}={\boldsymbol{v}}^{\intercal}{\boldsymbol{K}}_{\ell,\ell}{\boldsymbol{a}}\sim\|{\boldsymbol{K}}_{\ell,\ell}\|_{2}\|{\boldsymbol{a}}\|_{2}^{2}\sim\delta^{-d}_{\ell}\|{\boldsymbol{a}}\|_{2}^{2}.$
+Taking square roots yields the first claim. Similarly, we observe
+$\|v_{\ell}|X_{\ell}\|_{\ell^{2}}^{2}=\|{\boldsymbol{K}}_{\ell,\ell}{\boldsymbol{a}}\|_{2}^{2}={\boldsymbol{a}}^{\intercal}{\boldsymbol{K}}_{\ell,\ell}^{2}{\boldsymbol{a}}\sim\delta_{\ell}^{-2d}\|{\boldsymbol{a}}\|_{2}^{2}$
+and the second claim is again obtained by taking square roots. ∎
+Combining the previous two lemmata, we obtain a bound on the consistency error of the residual corrections.
+Theorem 4.3.
+Let $w_{\ell}=\sum_{i=1}^{N_{\ell}}c_{\ell,i}\Phi_{\ell}(\cdot-{\boldsymbol{x}}_{i})\in\mathcal{S}_{\ell}$ for $\ell=1,\ldots,L$ denote the residual corrections obtained from the solution of $\mathcal{K}{\boldsymbol{c}}={\boldsymbol{f}}$, see (10), and let $\widehat{w}_{\ell}=\sum_{i=1}^{N_{\ell}}\widehat{c}_{\ell,i}\Phi_{\ell}(\cdot-{\boldsymbol{x}}_{i})\in\mathcal{S}_{\ell}$ for $\ell=1,\ldots,L$ be the residual corrections obtained from solving the perturbed system $\widehat{\mathcal{K}}\widehat{\boldsymbol{c}}={\boldsymbol{f}}$.
+Then, there holds
+$\frac{\sum_{\ell=1}^{L}\|w_{\ell}-\widehat{w}_{\ell}\|_{\mathcal{N}_{\Phi_{\ell}}}^{2}}{\sum_{\ell=1}^{L}\|w_{\ell}\|_{\mathcal{N}_{\Phi_{\ell}}}^{2}}\leq C\frac{\big{\|}\mathcal{K}_{\mathrm{pc}}-\widehat{\mathcal{K}}_{\mathrm{pc}}\big{\|}_{F}^{2}}{\|\mathcal{K}_{\mathrm{pc}}\|_{F}^{2}}$
+for a constant $C>0$.
+Proof.
+Since the solutions to the original systems and their preconditioned versions are connected via ${\boldsymbol{d}}=\mathcal{D}^{-1/2}{\boldsymbol{c}}$ and $\widehat{\boldsymbol{d}}=\mathcal{D}^{-1/2}\widehat{\boldsymbol{c}}$, respectively, the first equivalence in Lemma 4.2 yields
+$\|w_{\ell}-\widehat{w}_{\ell}\|_{\mathcal{N}_{\Phi_{\ell}}}^{2}\sim\delta^{-d}_{\ell}\|{\boldsymbol{c}}_{\ell}-\widehat{\boldsymbol{c}}_{\ell}\|_{2}^{2}=\|{\boldsymbol{d}}_{\ell}-\widehat{\boldsymbol{d}}_{\ell}\|_{2}^{2}$
+as well as
+$\|w_{\ell}\|_{\mathcal{N}_{\Phi_{\ell}}}^{2}\sim\delta^{-d}_{\ell}\|{\boldsymbol{c}}_{\ell}\|_{2}^{2}=\|{\boldsymbol{d}}_{\ell}\|_{2}^{2}.$
+In view of $\|{\boldsymbol{d}}-\widehat{\boldsymbol{d}}\|_{2}^{2}=\sum_{\ell=1}^{L}\|{\boldsymbol{d}}_{\ell}-\widehat{\boldsymbol{d}}_{\ell}\|_{2}^{2}$ and $\|{\boldsymbol{d}}\|_{2}^{2}=\sum_{\ell=1}^{L}\|{\boldsymbol{d}}_{\ell}\|_{2}^{2}$, we arrive at
+$\frac{\sum_{\ell=1}^{L}\|w_{\ell}-\widehat{w}_{\ell}\|_{\mathcal{N}_{\Phi_{\ell}}}^{2}}{\sum_{\ell=1}^{L}\|w_{\ell}\|_{\mathcal{N}_{\Phi_{\ell}}}^{2}}\sim\frac{\|{\boldsymbol{d}}-\widehat{\boldsymbol{d}}\|_{2}^{2}}{\|{\boldsymbol{d}}\|_{2}^{2}}\leq C\frac{\big{\|}\mathcal{K}_{\mathrm{pc}}-\widehat{\mathcal{K}}_{\mathrm{pc}}\big{\|}_{F}^{2}}{\|\mathcal{K}_{\mathrm{pc}}\|_{F}^{2}}$
+by employing Lemma 4.1. This completes the proof. ∎
+To bound the numerical approximation error between $s_{L}$ and $\widehat{s}_{L}\mathrel{\mathrel{\mathop{:}}=}\widehat{w}_{1}+\dots+\widehat{w}_{L}$, we have the following result, which links this error to the approximation error of the residual corrections.
+Lemma 4.4.
+Under the conditions of Theorem 4.3, the approximation $\widehat{s}_{L}\mathrel{\mathrel{\mathop{:}}=}\widehat{w}_{1}+\dots+\widehat{w}_{L}$ to $s_{L}$ satisfies the estimate
+$\|s_{L}-\widehat{s}_{L}\|_{L^{2}(\Omega)}^{2}\leq CL\sum_{\ell=1}^{L}\|w_{\ell}-\widehat{w}_{\ell}\|_{\mathcal{N}_{\Phi_{\ell}}}^{2}.$
+for some constant $C>0$.
+Proof.
+From the Cauchy–Schwarz inequality, it follows that
+$\|s_{L}-\widehat{s}_{L}\|^{2}_{L^{2}(\Omega)}=\bigg{\|}\sum_{\ell=1}^{L}(w_{\ell}-\widehat{w}_{\ell})\bigg{\|}^{2}_{L^{2}(\Omega)}\leq L\sum_{\ell=1}^{L}\|w_{\ell}-\widehat{w}_{\ell}\|^{2}_{L^{2}(\Omega)}.$
+Next, we consider the sampling inequality from [16, Theorem 3.5], i.e.,
+$\|u\|_{L^{p}}\leq C\Big{(}h_{X,\Omega}^{\theta}\|u\|_{H^{\theta}(\Omega)}+h_{X,\Omega}^{d/p}\|u|X\|_{\ell^{2}}\Big{)},\quad C>0.$
+Letting $p=2$ and squaring the previous inequality results in
+$\|u\|_{L^{2}(\Omega)}^{2}\leq 2C\Big{(}h_{X,\Omega}^{2\theta}\|u\|_{H^{\theta}(\Omega)}^{2}+h_{X,\Omega}^{d}\|u|X\|_{\ell^{2}}^{2}\Big{)}.$
+Inserting this in the first inequality, we arrive at
+$\|s_{L}-\widehat{s}_{L}\|^{2}_{L^{2}(\Omega)}\leq 2CL\sum_{\ell=1}^{L}\big{(}h_{\ell}^{2\theta}\|w_{\ell}-\widehat{w}_{\ell}\|_{H^{\theta}(\Omega)}^{2}+h_{\ell}^{d}\|w_{\ell}-\widehat{w}_{\ell}|X_{\ell}\|_{\ell^{2}}^{2}\big{)}.$
+Next, we have from [29, Lemma 1] that
+$\|w_{\ell}-\widehat{w}_{\ell}\|^{2}_{H^{\theta}(\Omega)}\leq C\delta_{\ell}^{-2\theta}\|w_{\ell}-\widehat{w}_{\ell}\|^{2}_{\mathcal{N}_{\Phi_{\ell}}},\quad C>0,$
+while the second equivalence in Lemma 4.2 amounts to
+$\|w_{\ell}-\widehat{w}_{\ell}|X_{\ell}\|_{\ell^{2}}^{2}\sim\delta_{\ell}^{-2d}\|{\boldsymbol{c}}_{\ell}-\widehat{\boldsymbol{c}}_{\ell}\|_{2}^{2}\sim\delta_{\ell}^{-d}\|w_{\ell}-\widehat{w}_{\ell}\|^{2}_{\mathcal{N}_{\Phi_{\ell}}}.$
+Inserting these two estimates in the previous one yields with $\delta_{\ell}=\eta h_{\ell}$, cp. (4), that
+$\|s_{L}-\widehat{s}_{L}\|^{2}_{L^{2}(\Omega)}\leq CL\sum_{\ell=1}^{L}\big{(}\eta^{-2\theta}\|w_{\ell}-\widehat{w}_{\ell}\|^{2}_{\mathcal{N}_{\Phi_{\ell}}}+\eta^{-d}\|w_{\ell}-\widehat{w}_{\ell}\|^{2}_{\mathcal{N}_{\Phi_{\ell}}}\big{)},$
+for some constant $C>0$, from which the claim is immediately obtained. ∎
+Finally, we need a bound on the residual corrections by the right hand side $f$.
+Lemma 4.5.
+Under the conditions of Theorem 2.1, there exists a constant $C>0$ such that
+$\sum_{\ell=1}^{L}\|w_{\ell}\|_{\mathcal{N}_{\Phi_{\ell}}}^{2}\leq C\|f\|^{2}_{H^{\theta}(\Omega)}.$
+Proof.
+By construction, the residual correction $w_{\ell}$ is the norm-minimal interpolant to the residual $e_{\ell-1}\mathrel{\mathrel{\mathop{:}}=}f-\sum_{m=1}^{\ell-1}w_{m}$, for any $1\leq\ell\leq L$, which implies
+$\|w_{\ell}\|_{\mathcal{N}_{\Phi_{\ell}}}\leq\|Ee_{\ell-1}||_{\mathcal{N}_{\Phi_{\ell}}}$
+with the continuous extension operator $E\colon H^{\theta}(\Omega)\to H^{\theta}(\mathbb{R}^{d})$, $\|Ef\|_{H^{\theta}(\mathbb{R}^{d})}\leq C_{E}\|f\|_{H^{\theta}(\Omega)}$.
+Now we can use the recursion estimate [29, Theorem 1] to obtain
+$\|Ee_{\ell-1}||_{\mathcal{N}_{\Phi_{\ell}}}\leq(C_{1}\mu^{\theta}+C_{1}\eta^{-\theta})\|Ee_{\ell-2}\|_{\mathcal{N}_{\Phi_{\ell-1}}}.$
+Recurring $\ell-2$ times, we arrive at the bound
+$\|w_{\ell}\|_{\mathcal{N}_{\Phi_{\ell}}}\leq\|Ee_{\ell-1}||_{\mathcal{N}_{\Phi_{\ell}}}$ $\leq(C_{1}\mu^{\theta}+C_{1}\eta^{-\theta})^{\ell-2}\|Ee_{0}\|_{\mathcal{N}_{\Phi_{1}}}$
+$=(C_{1}\mu^{\theta}+C_{1}\eta^{-\theta})^{\ell-2}\|Ef\|_{\mathcal{N}_{\Phi_{1}}}$
+$\leq C_{E}(C_{1}\mu^{\theta}+C_{1}\eta^{-\theta})^{\ell-2}\delta_{1}^{-\theta}\|f\|_{H^{\theta}(\Omega)}.$
+By this bound and by employing the summation formula for the geometric series, we finally estimate
+$\sum_{\ell=1}^{L}\|w_{\ell}\|_{\mathcal{N}_{\Phi_{\ell}}}^{2}$ $\leq\sum_{\ell=1}^{L}C_{E}^{2}c(C_{1}\mu^{\theta}+C_{1}\eta^{-\theta})^{2(\ell-2)}\delta_{1}^{-2\theta}\|f\|_{H^{\theta}(\Omega)}^{2}$
+$=\frac{C_{E}^{2}\delta_{1}^{-2\theta}}{(C_{1}\mu^{\theta}+C_{1}\eta^{-\theta})^{2}}\|f\|_{H^{\theta}(\Omega)}^{2}\sum_{\ell=0}^{L-1}(C_{1}\mu^{\theta}+C_{1}\eta^{-\theta})^{2\ell}$
+$=\frac{C_{E}^{2}\delta_{1}^{-2\theta}}{(C_{1}\mu^{\theta}+C_{1}\eta^{-\theta})^{2}}\frac{1-(C_{1}\mu^{\theta}+C_{1}\eta^{-\theta})^{2L}}{1-(C_{1}\mu^{\theta}+C_{1}\eta^{-\theta})^{2}}\|f\|_{H^{\theta}(\Omega)}^{2}.$
+This completes the proof. ∎
+Combining the previous lemmata with Theorem 2.1, we arrive at the final error estimate.
+Theorem 4.6.
+Under the conditions of Theorems 2.1 and 4.3, there exists a constant $C>0$ such that
+$\|f-\widehat{s}_{L}\|_{L^{2}(\Omega)}\leq C\Bigg{(}(C_{1}\mu^{\theta}+C_{1}\eta^{-\theta})^{L}+\sqrt{L}\frac{\big{\|}\mathcal{K}_{\mathrm{pc}}-\widehat{\mathcal{K}}_{\mathrm{pc}}\big{\|}_{F}}{\|\mathcal{K}_{\mathrm{pc}}\|_{F}}\Bigg{)}\|f\|_{H^{\theta}(\Omega)}.$
+Proof.
+By the triangle inequality, there holds
+$\|f-\widehat{s}_{L}\|_{L^{2}(\Omega)}\leq\|f-{s}_{L}\|_{L^{2}(\Omega)}+\|s_{L}-\widehat{s}_{L}\|_{L^{2}(\Omega)}.$
+The first term on the right hand side is bounded using Theorem 2.1. For the second term, we obtain using Lemmata 4.4 and 4.5 as well as Theorem 4.3 that
+$\|s_{L}-\widehat{s}_{L}\|_{L^{2}(\Omega)}^{2}$ $\leq CL\sum_{\ell=1}^{L}\|w_{\ell}-\widehat{w}_{\ell}\|^{2}_{\mathcal{N}_{\Phi_{\ell}}}\leq CL\frac{\big{\|}\mathcal{K}_{\mathrm{pc}}-\widehat{\mathcal{K}}_{\mathrm{pc}}\big{\|}_{F}^{2}}{\|\mathcal{K}_{\mathrm{pc}}\|_{F}^{2}}\sum_{\ell=1}^{L}\|w_{\ell}\|^{2}_{\mathcal{N}_{\Phi_{\ell}}}$
+$\leq CL\frac{\big{\|}\mathcal{K}_{\mathrm{pc}}-\widehat{\mathcal{K}}_{\mathrm{pc}}\big{\|}_{F}^{2}}{\|\mathcal{K}_{\mathrm{pc}}\|_{F}^{2}}\|f\|_{H^{\theta}(\Omega)}^{2}.$
+Taking square roots and inserting into the first inequality yields the claim. ∎
 5. Samplets for scattered data approximation
@@
 We briefly recall their underlying concepts as introduced in [9].
-To this end, we consider the image of the spaces $\mathcal{S}_{\ell}$, $\ell=1,\ldots,L$, under the Riesz isomorphism $\mathcal{J}\colon\mathcal{N}_{\Phi_{\ell}}\to\mathcal{N}^{\prime}_{\Phi_{\ell}}$ and define
+To this end, we consider the image of the spaces $\mathcal{S}_{\ell}$, $\ell=1,\ldots,L$, defined in Equation (7), under the Riesz isomorphism $\mathcal{J}\colon\mathcal{N}_{\Phi_{\ell}}\to\mathcal{N}^{\prime}_{\Phi_{\ell}}$ and define
 (14) $\mathcal{S}_{\ell}^{\prime}\mathrel{\mathrel{\mathop{:}}=}\operatorname{span}\{\mathcal{J}\Phi_{\ell}(\cdot-{\boldsymbol{x}}):{\boldsymbol{x}}\in X_{\ell}\}=\operatorname{span}\{\delta_{\boldsymbol{x}}:{\boldsymbol{x}}\in X_{\ell}\},$
 by the identity $\mathcal{J}\Phi_{\ell}(\cdot-{\boldsymbol{x}})=\delta_{\boldsymbol{x}}$, where $\delta_{\boldsymbol{x}}$ is the point evaluation functional at ${\boldsymbol{x}}\in\Omega$.
-On the spaces $\mathcal{S}_{\ell}^{\prime}$, we introduce an inner product different from the canonical one by the defining property $\langle\delta_{{\boldsymbol{x}}_{i}},\delta_{{\boldsymbol{x}}_{j}}\rangle_{\mathcal{S}_{\ell}^{\prime}}\mathrel{\mathrel{\mathop{:}}=}\delta_{ij}$ for ${\boldsymbol{x}}_{i},{\boldsymbol{x}}_{j}\in X_{\ell}$.
+We equip the spaces $\mathcal{S}_{\ell}^{\prime}$ with an inner product, different from the canonical one, by the defining property $\langle\delta_{{\boldsymbol{x}}_{i}},\delta_{{\boldsymbol{x}}_{j}}\rangle_{\mathcal{S}_{\ell}^{\prime}}\mathrel{\mathrel{\mathop{:}}=}\delta_{ij}$ for ${\boldsymbol{x}}_{i},{\boldsymbol{x}}_{j}\in X_{\ell}$.
 For the sake of a more lightweight notation, we drop the subscript $\ell$ of the spaces $\mathcal{S}_{\ell}^{\prime}$ for the remainder of this paragraph and remark that each space $\mathcal{S}_{\ell}^{\prime}$ will in general have a different multiresolution analysis, where the maximum level $J$ depends on $\ell$.
-Given a multiresolution analysis
-(14) $\mathcal{V}_{0}^{\prime}\subset\mathcal{V}_{1}^{\prime}\subset\cdots\subset\mathcal{V}_{J}^{\prime}=\mathcal{S}^{\prime},$
-we keep track of the increment of information between two consecutive levels $j$ and $j+1$. Since there holds $\mathcal{V}_{j}^{\prime}\subset\mathcal{V}_{j+1}^{\prime}$, we may orthogonally decompose
-$\mathcal{V}_{j+1}^{\prime}=\mathcal{V}_{j}^{\prime}\overset{\perp}{\oplus}\mathcal{W}_{j}^{\prime}$
-for a certain detail space $\mathcal{W}_{j}^{\prime}$.
+Given a multiresolution analysis $\mathcal{V}_{0}^{\prime}\subset\mathcal{V}_{1}^{\prime}\subset\cdots\subset\mathcal{V}_{J}^{\prime}=\mathcal{S}^{\prime},$ we keep track of the increment of information between two consecutive levels $j$ and $j+1$.
+Since there holds $\mathcal{V}_{j}^{\prime}\subset\mathcal{V}_{j+1}^{\prime}$, we may orthogonally decompose $\mathcal{V}_{j+1}^{\prime}=\mathcal{V}_{j}^{\prime}{\oplus}\mathcal{W}_{j}^{\prime}$ for a certain detail space $\mathcal{W}_{j}^{\prime}\perp\mathcal{V}_{j}^{\prime}$.
 In analogy to wavelet nomenclature, we call the elements of a basis of $\mathcal{V}_{0}^{\prime}$ scaling distributions and the elements of a basis of one of the spaces $\mathcal{W}_{j}^{\prime}$ samplets.
@@
 A samplet basis can be constructed such that it exhibits vanishing moments of order $q+1$, i.e.,
-(15) $(p,\sigma_{j,k})_{\Omega}=0\ \text{for all}\ p\in\mathcal{P}_{q},$
+(15) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\sigma_{j,k}(p)=0\quad\text{for all}\ p\in\mathcal{P}_{q},}$
 where $\sigma_{j,k}\in\mathcal{W}_{j}^{\prime}$ is a samplet and $\mathcal{P}_{q}\mathrel{\mathrel{\mathop{:}}=}\operatorname{span}\{{\boldsymbol{x}}^{\boldsymbol{\alpha}}:\|\boldsymbol{\alpha}\|_{1}\leq q\}$ denotes the space of polynomials of total degree at most $q$.
@@
 Such a cluster tree $\mathcal{C}$ directly induces a support based hierarchical clustering of the subspace $\mathcal{S}^{\prime}$, spanned by the Dirac-$\delta$-distributions supported at the data sites in $X$.
-With a slight abuse of notation, we will refer to this cluster tree also by $\mathcal{C}$ and to its elements by $\tau$.
+With a slight abuse of notation, we will refer to this cluster tree also by $\mathcal{C}$ and to its nodes by $\tau$.
 Given a cluster tree $\mathcal{C}$, a samplet basis for $\mathcal{S}^{\prime}$, where $\dim\mathcal{S}^{\prime}=N$, can be constructed with cost $\mathcal{O}(N)$.
@@
   3. (3)
-    The coefficient vector ${\boldsymbol{\omega}}_{j,k}=\big{[}\omega_{j,k,i}\big{]}_{i}$ of the samplet $\sigma_{j,k}$ satisfies the bound
-    $\|{\boldsymbol{\omega}}_{j,k}\|_{1}\leq c2^{(J-j)/2}$
-    for a constant $20<c<$ 2.
+    The coefficient vector ${\boldsymbol{\omega}}_{j,k}=\big{[}\omega_{j,k,i}\big{]}_{i}$ of the samplet $\sigma_{j,k}$ satisfies the bound $\|{\boldsymbol{\omega}}_{j,k}\|_{1}\leq c2^{(J-j)/2}$ for a constant $20<c<$ 2.
   4. (4)
@@
     For quasi-uniform sets $X$, there holds $\operatorname{diam}\operatorname{supp}\sigma_{j,k}\leq c2^{-j/d}$ for some constant $c>0$.
-A direct consequence of the orthonormality of the samplet basis is that the samplet transform
-$[\sigma_{j,k}]_{j,k}={\boldsymbol{T}}[\delta_{{\boldsymbol{x}}_{i}}]_{i=1}^{N}$
-satisfies ${\boldsymbol{T}}^{\intercal}{\boldsymbol{T}}={\boldsymbol{T}}{\boldsymbol{T}}^{\intercal}={\boldsymbol{I}}\in\mathbb{R}^{N\times N}$.
+A direct consequence of the orthonormality of the samplet basis is that the samplet transform $[\sigma_{j,k}]_{j,k}={\boldsymbol{T}}[\delta_{{\boldsymbol{x}}_{i}}]_{i=1}^{N}$ satisfies ${\boldsymbol{T}}^{\intercal}{\boldsymbol{T}}={\boldsymbol{T}}{\boldsymbol{T}}^{\intercal}={\boldsymbol{I}}\in\mathbb{R}^{N\times N}$.
 If the samplet transform ${\boldsymbol{T}}$ is computed recursively adhering the hierarchy induced by the cluster tree $\mathcal{C}$, its cost is of order $\mathcal{O}(N)$.
-Employing the Riesz isometry, the samplet basis gives rise to a basis for $\mathcal{S}\subset\mathcal{N}_{\Phi}$. To this end, consider a samplet
-$\sigma_{j,k}=\sum_{i=1}^{N}\omega_{j,k,i}\delta_{{\boldsymbol{x}}_{i}}.$
-The samplet can be identified with the function
-$\psi_{j,k}\mathrel{\mathrel{\mathop{:}}=}\sum_{i=1}^{N}\omega_{j,k,i}\Phi({\boldsymbol{x}}_{i}-\cdot)\in\mathcal{S}$
-by means of the Riesz isometry. The vanishing moment property (15) then translates to
-$\langle\psi_{j,k},f\rangle_{\mathcal{N}_{\Phi}}=0$
-for any $f\in\mathcal{N}_{\Phi}$ which satisfies $f|_{O}\in\mathcal{P}_{q}$ for any open and convex set $O$ with $\operatorname{supp}(\sigma_{j,k})\subset O\cap\Omega$.
+Employing the Riesz isometry, the samplet basis induces a basis for $\mathcal{S}\subset\mathcal{N}_{\Phi}$.
+To this end, consider a samplet $\sigma_{j,k}=\sum_{i=1}^{N}\omega_{j,k,i}\delta_{{\boldsymbol{x}}_{i}}.$ The samplet can be identified with the function $\psi_{j,k}\mathrel{\mathrel{\mathop{:}}=}\sum_{i=1}^{N}\omega_{j,k,i}\Phi({\boldsymbol{x}}_{i}-\cdot)\in\mathcal{S}$ by means of the Riesz isometry.
+The vanishing moment property (15) then translates to $\langle\psi_{j,k},f\rangle_{\mathcal{N}_{\Phi}}=0$ for any $f\in\mathcal{N}_{\Phi}$ which satisfies $f|_{\operatorname{supp}(\sigma_{j,k})}\in\mathcal{P}_{q}$.
 Furthermore, there holds
@@
 This means that the Gramian of the embedded samplet basis in $\mathcal{S}$ is just the samplet transformed generalized Vandermonde matrix.
-For asymptotically smooth kernels $K$, i.e.
-that there exist $C,r>0$ such that for all $({\boldsymbol{x}},{\boldsymbol{y}})\in(\Omega\times\Omega)\setminus\Delta$
+For asymptotically smooth kernels $K$, i.e., there exist $C,r>0$ such that for all $({\boldsymbol{x}},{\boldsymbol{y}})\in(\Omega\times\Omega)\setminus\Delta$
 (17) $\bigg{|}\frac{\partial^{|\boldsymbol{\alpha}|+|\boldsymbol{\beta}|}}{\partial{\boldsymbol{x}}^{\boldsymbol{\alpha}}\partial{\boldsymbol{y}}^{\boldsymbol{\beta}}}K({\boldsymbol{x}},{\boldsymbol{y}})\bigg{|}\leq C\frac{(|\boldsymbol{\alpha}|+|\boldsymbol{\beta}|)!}{r^{|\boldsymbol{\alpha}|+|\boldsymbol{\beta}|}\|{\boldsymbol{x}}-{\boldsymbol{y}}\|_{2}^{|\boldsymbol{\alpha}|+|\boldsymbol{\beta}|}}$
 uniformly in $\boldsymbol{\alpha},\boldsymbol{\beta}\in\mathbb{N}^{d}$ apart from the diagonal $\Delta\mathrel{\mathrel{\mathop{:}}=}\{({\boldsymbol{x}},{\boldsymbol{y}})\in\Omega\times\Omega:{\boldsymbol{x}}={\boldsymbol{y}}\}$, the matrix ${\boldsymbol{K}}^{\Sigma}$ becomes quasi-sparse and can efficiently be computed.
-More precisely, there holds the following result, see [7].
+More precisely, a compression based on cluster distances is applied to compute a sparsified version of the matrix ${\boldsymbol{K}}^{\Sigma}$.
+In detail, there holds the following result, see [9].
 Theorem 5.2.
@@
 (18) $\operatorname{dist}(\tau,\tau^{\prime})\geq\rho\max\{\operatorname{diam}(\tau),\operatorname{diam}(\tau^{\prime})\},\quad\rho>0,$
-where $\tau$ is the cluster supporting $\sigma_{j,k}$ and $\tau^{\prime}$ is the cluster supporting $\sigma_{j^{\prime},k^{\prime}}$, respectively. Then, there exists a constant $C>0$, such that
-(19) $\frac{\big{\|}{\boldsymbol{K}}^{\Sigma}-{\boldsymbol{K}}^{\Sigma}_{\varepsilon}\big{\|}_{F}}{\big{\|}{\boldsymbol{K}}^{\Sigma}\big{\|}_{F}}\leq Cm_{q}(r\rho/d)^{-2(q+1)},$
-where $m_{q}=\binom{q+d}{d}$ is the dimension of $\mathcal{P}_{q}$ , and the compressed matrix has $\mathcal{O}(m_{q}^{2}N\log N)$ nonzero coefficients.
+where $\tau$ is the cluster supporting $\sigma_{j,k}$ and $\tau^{\prime}$ is the cluster supporting $\sigma_{j^{\prime},k^{\prime}}$, respectively.
+Then, there exists a constant $C>0$, such that the resulting compressed matrix ${\boldsymbol{K}}^{\Sigma,\rho}$ satisfies
+(19) $\frac{\big{\|}{\boldsymbol{K}}^{\Sigma}-{\boldsymbol{K}}^{\Sigma,\rho}\big{\|}_{F}}{\big{\|}{\boldsymbol{K}}^{\Sigma}\big{\|}_{F}}\leq Cm_{q}\bigg{(}\frac{r\rho}{d}\bigg{)}^{-2(q+1)},$
+where $m_{q}=\binom{q+d}{d}$ is the dimension of $\mathcal{P}_{q}$. The compressed matrix has $\mathcal{O}(m_{q}^{2}N\log N)$ nonzero coefficients.
+We remark that, given a radial function satisfying (17), the corresponding scaled kernel (6) satisfies the asymptotical smoothness estimate, cp. [10],
+$\bigg{|}\frac{\partial^{|\boldsymbol{\alpha}|+|\boldsymbol{\beta}|}}{\partial{\boldsymbol{x}}^{\boldsymbol{\alpha}}\partial{\boldsymbol{y}}^{\boldsymbol{\beta}}}K_{\ell}({\boldsymbol{x}},{\boldsymbol{y}})\bigg{|}\leq C\delta_{\ell}^{-d}\frac{(|\boldsymbol{\alpha}|+|\boldsymbol{\beta}|)!}{(r\delta_{\ell})^{|\boldsymbol{\alpha}|+|\boldsymbol{\beta}|}\|({\boldsymbol{x}}-{\boldsymbol{y}})/\delta_{\ell}\|_{2}^{|\boldsymbol{\alpha}|+|\boldsymbol{\beta}|}}=C\delta_{\ell}^{-d}\frac{(|\boldsymbol{\alpha}|+|\boldsymbol{\beta}|)!}{r^{|\boldsymbol{\alpha}|+|\boldsymbol{\beta}|}\|{\boldsymbol{x}}-{\boldsymbol{y}}\|_{2}^{|\boldsymbol{\alpha}|+|\boldsymbol{\beta}|}},$
+which means that the dependency of the derivatives on $\delta_{\ell}$ just cancels out.
+Moreover, we remark that, since we only require derivatives of the kernel up to order $12q+$ 1 to be bounded, the asymptotical smoothness requirement could in principle also be weakened to kernels of finite smoothness, see, e.g., [11] for details.
 6. Multiscale algorithm in samplet coordinates
-In this section, we transform the interpolation problem from the original basis, referred to as the natural basis, to the samplet basis, which consists of the scaling distributions and the samplets, as introduced in Section 4.
+In this section, we transform the interpolation problem from the original basis, referred to as the kernel basis, to the samplet basis, as introduced in Section 5.
 In what follows, we denote the samplet transform for $\mathcal{S}^{\prime}_{\ell}$ by ${\boldsymbol{T}}_{\ell}\in\mathbb{R}^{N_{\ell}\times N_{\ell}}$ and recall that ${\boldsymbol{T}}^{\intercal}_{\ell}{\boldsymbol{T}}_{\ell}={\boldsymbol{T}}_{\ell}{\boldsymbol{T}}^{\intercal}_{\ell}={\boldsymbol{I}}\in\mathbb{R}^{N_{\ell}\times N_{\ell}}$.
-Defining the block transform
-(20) $\mathcal{T}\mathrel{\mathrel{\mathop{:}}=}\left[\begin{array}[]{cccc}{\boldsymbol{T}}_{1}&&&\\ &{\boldsymbol{T}}_{2}&&\\ &&\ddots&\\ &&&{\boldsymbol{T}}_{L}\end{array}\right],$
-there holds $\mathcal{T}^{\intercal}\mathcal{T}=\mathcal{T}\mathcal{T}^{\intercal}={\boldsymbol{I}}$. Hence, the linear system (8) is equivalent to
-$\mathcal{T}^{\intercal}\mathcal{T}\left[\begin{array}[]{cccc}\boldsymbol{K}_{1,1}&&&\\ \boldsymbol{K}_{2,1}&\boldsymbol{K}_{2,2}&&\\ \vdots&\cdots&\ddots&\\ \boldsymbol{K}_{L,1}&\cdots&\boldsymbol{K}_{L,L-1}&\boldsymbol{K}_{L,L}\end{array}\right]\mathcal{T}^{\intercal}\mathcal{T}\left[\begin{array}[]{c}\boldsymbol{c}_{1}\\ \boldsymbol{c}_{2}\\ \boldsymbol{c}_{3}\\ \vdots\\ \boldsymbol{c}_{L}\end{array}\right]=\mathcal{T}^{\intercal}\mathcal{T}\left[\begin{array}[]{c}\boldsymbol{f}_{1}\\ \boldsymbol{f}_{2}\\ \boldsymbol{f}_{3}\\ \vdots\\ \boldsymbol{f}_{L}\end{array}\right]$
-Letting $\boldsymbol{K}^{\Sigma}_{\ell,\ell^{\prime}}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{T}}_{\ell}\boldsymbol{K}_{\ell,\ell^{\prime}}{\boldsymbol{T}}^{\intercal}_{\ell^{\prime}}$, $\boldsymbol{c}^{\Sigma}_{\ell}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{T}}_{\ell}\boldsymbol{c}_{\ell}$ and $\boldsymbol{f}^{\Sigma}_{\ell}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{T}}_{\ell}\boldsymbol{f}_{\ell}$ and dropping the inverse block transform, we arrive at the samplet transformed system
-(21) $\left[\begin{array}[]{ccccc}\boldsymbol{K}^{\Sigma}_{1,1}&&&&\\ \boldsymbol{K}^{\Sigma}_{2,1}&\boldsymbol{K}^{\Sigma}_{2,2}&&&\\ \boldsymbol{K}^{\Sigma,3}_{1}&\boldsymbol{K}^{\Sigma}_{3,2}&\boldsymbol{K}^{\Sigma}_{3,3}&&\\ \vdots&\vdots&\cdots&\ddots&\\ \boldsymbol{K}^{\Sigma}_{L,1}&\boldsymbol{K}^{\Sigma}_{L,2}&\cdots&\boldsymbol{K}^{\Sigma}_{L,L-1}&\boldsymbol{K}^{\Sigma}_{L,L}\end{array}\right]\left[\begin{array}[]{c}\boldsymbol{c}^{\Sigma}_{1}\\ \boldsymbol{c}^{\Sigma}_{2}\\ \boldsymbol{c}^{\Sigma}_{3}\\ \vdots\\ \boldsymbol{c}^{\Sigma}_{L}\end{array}\right]=\left[\begin{array}[]{c}\boldsymbol{f}^{\Sigma}_{1}\\ \boldsymbol{f}^{\Sigma}_{2}\\ \boldsymbol{f}^{\Sigma}_{3}\\ \vdots\\ \boldsymbol{f}^{\Sigma}_{L}\end{array}\right],$
-which is equivalent to (8). In particular, since $\mathcal{T}$ is an isometry, the condition number of the system (21) is identical to that of (8).
+Defining the block transform $\mathcal{T}\mathrel{\mathrel{\mathop{:}}=}\operatorname{diag}({\boldsymbol{T}}_{1},\ldots,{\boldsymbol{T}}_{L})$ there holds $\mathcal{T}^{\intercal}\mathcal{T}=\mathcal{T}\mathcal{T}^{\intercal}={\boldsymbol{I}}$.
+Hence, the linear system (10) is equivalent to
+$\mathcal{T}^{\intercal}\mathcal{T}\left[\begin{array}[]{cccc}\boldsymbol{K}_{1,1}&&&\\ \boldsymbol{K}_{2,1}&\boldsymbol{K}_{2,2}&&\\ \vdots&\vdots&\ddots&\\ \boldsymbol{K}_{L,1}&\boldsymbol{K}_{L,2}&\cdots&\boldsymbol{K}_{L,L}\end{array}\right]\mathcal{T}^{\intercal}\mathcal{T}\left[\begin{array}[]{c}\boldsymbol{c}_{1}\\ \boldsymbol{c}_{2}\\ \vdots\\ \boldsymbol{c}_{L}\end{array}\right]=\mathcal{T}^{\intercal}\mathcal{T}\left[\begin{array}[]{c}\boldsymbol{f}_{1}\\ \boldsymbol{f}_{2}\\ \vdots\\ \boldsymbol{f}_{L}\end{array}\right].$
+Letting $\boldsymbol{K}^{\Sigma}_{\ell,\ell^{\prime}}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{T}}_{\ell}\boldsymbol{K}_{\ell,\ell^{\prime}}{\boldsymbol{T}}^{\intercal}_{\ell^{\prime}}\in\mathbb{R}^{N_{\ell}\times N_{\ell}^{\prime}},\ \boldsymbol{c}^{\Sigma}_{\ell}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{T}}_{\ell}\boldsymbol{c}_{\ell}\in\mathbb{R}^{N_{\ell}},\ \boldsymbol{f}^{\Sigma}_{\ell}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{T}}_{\ell}\boldsymbol{f}_{\ell}\in\mathbb{R}^{N_{\ell}},$ we can rewrite the previous system according to
+(20) $\left[\begin{array}[]{cccc}\boldsymbol{K}^{\Sigma}_{1,1}&&&\\ \boldsymbol{K}^{\Sigma}_{2,1}&\boldsymbol{K}^{\Sigma}_{2,2}&&\\ \vdots&\vdots&\ddots&\\ \boldsymbol{K}^{\Sigma}_{L,1}&\boldsymbol{K}^{\Sigma}_{L,2}&\cdots&\boldsymbol{K}^{\Sigma}_{L,L}\end{array}\right]\left[\begin{array}[]{c}\boldsymbol{c}^{\Sigma}_{1}\\ \boldsymbol{c}^{\Sigma}_{2}\\ \vdots\\ \boldsymbol{c}^{\Sigma}_{L}\end{array}\right]=\left[\begin{array}[]{c}\boldsymbol{f}^{\Sigma}_{1}\\ \boldsymbol{f}^{\Sigma}_{2}\\ \vdots\\ \boldsymbol{f}^{\Sigma}_{L}\end{array}\right].$
+Since $\mathcal{T}$ is an isometry, the condition number of this system is identical to that of (10).
+In particular, the diagonal blocks remain well-conditioned in accordance with Corollary 3.2.
+Applying the matrix compression from Theorem 5.2 to each of the blocks $\boldsymbol{K}^{\Sigma}_{\ell,\ell^{\prime}}$, we arrive at the compressed, samplet transformed system
+(21) $\left[\begin{array}[]{cccc}\boldsymbol{K}^{\Sigma,\rho}_{1,1}&&&\\ \boldsymbol{K}^{\Sigma,\rho}_{2,1}&\boldsymbol{K}^{\Sigma,\rho}_{2,2}&&\\ \vdots&\vdots&\ddots&\\ \boldsymbol{K}^{\Sigma,\rho}_{L,1}&\boldsymbol{K}^{\Sigma,\rho}_{L,2}&\cdots&\boldsymbol{K}^{\Sigma,\rho}_{L,L}\end{array}\right]\left[\begin{array}[]{c}\boldsymbol{c}^{\Sigma}_{1}\\ \boldsymbol{c}^{\Sigma}_{2}\\ \vdots\\ \boldsymbol{c}^{\Sigma}_{L}\end{array}\right]=\left[\begin{array}[]{c}\boldsymbol{f}^{\Sigma}_{1}\\ \boldsymbol{f}^{\Sigma}_{2}\\ \vdots\\ \boldsymbol{f}^{\Sigma}_{L}\end{array}\right].$
+A detailed algorithm for solving (21) using block forward substitution and evaluating the resulting approximation at a set of points can be found in Algorithm (1).
+We remark that replacing the kernel matrix for the evaluation points by its samplet compressed version leads to a further consistency error, which can, however, be bounded with similar ideas as in Lemma 4.1.
+Algorithm 1 Multiscale interpolation in samplet coordinates
+Input: Sets of points $\{X_{1},\ldots,X_{L}\}$, function $f$, parameters $\rho,q$.
+Output: Coefficient vectors $\widehat{\boldsymbol{c}}_{1},\ldots,\widehat{\boldsymbol{c}}_{L}$ corresponding to $\widehat{w}_{1},\ldots,\widehat{w}_{L}$.
+1:  for $\ell=1,2,\ldots,L$ do
+2:     Compute ${\boldsymbol{f}}^{\Sigma}_{\ell}\leftarrow{\boldsymbol{T}}_{\ell}{\boldsymbol{f}}_{\ell}\mathrel{\mathrel{\mathop{:}}=}f|X_{\ell}$.
+3:     for $\ell^{\prime}=1,2,\ldots,\ell$ do
+4:        Compute $\boldsymbol{K}^{\Sigma,\rho}_{\ell,\ell^{\prime}}$ using $q+1$ vanishing moments.
+5:     end for
+6:     Compute $\widehat{\boldsymbol{e}}^{\Sigma}_{\ell}\leftarrow\boldsymbol{f}^{\Sigma}_{\ell}-\sum_{\ell^{\prime}=1}^{\ell-1}\boldsymbol{K}^{\Sigma,\rho}_{\ell,\ell^{\prime}}\widehat{\boldsymbol{c}}^{\Sigma}_{\ell^{\prime}}$.
+7:     Solve the local interpolation problem $\boldsymbol{K}^{\Sigma,\rho}_{\ell,\ell}\widehat{\boldsymbol{c}}^{\Sigma}_{\ell}=\widehat{\boldsymbol{e}}^{\Sigma}_{\ell}$.
+8:     Compute $\widehat{\boldsymbol{c}}_{\ell}\leftarrow{\boldsymbol{T}}_{\ell}^{\intercal}\widehat{\boldsymbol{c}}^{\Sigma}_{\ell}$.
+9:  end for
+The overall compression error of (21), can be estimated by summing up the block-wise errors in accordance with Theorem 5.2. The following result is immediate.
+Theorem 6.1.
+Let $X_{1},\ldots,X_{L}$ be quasi-uniform. Then, the compression error in the linear system (21) can be bounded according to
+$\frac{\big{\|}{\mathcal{K}}^{\Sigma}-{\mathcal{K}}^{\Sigma,\rho}\big{\|}_{F}}{\big{\|}{\mathcal{K}}^{\Sigma}\big{\|}_{F}}\leq Cm_{q}(r\rho/d)^{-2(q+1)},$
+where we denote the block matrix from (20) by ${\mathcal{K}}^{\Sigma}$ and the one from (21) by ${\mathcal{K}}^{\Sigma,\rho}$, respectively.
+Observing that $\mathcal{T}\mathcal{D}^{1/2}=\mathcal{D}^{1/2}\mathcal{T}$, combining the previous theorem with Theorem 4.6 immediately results in an error estimate for the approximation $\widehat{s}_{L}$ to $s_{L}$ obtained by solving the samplet compressed system.
+Theorem 6.2.
+Let the conditions of Theorem 4.6 be satisfied. Then, the numerical approximation $\widehat{s}_{L}$ resulting from solving (21) satisfies the error estimate
+$\|f-\widehat{s}_{L}\|_{L^{2}(\Omega)}\leq C\Big{(}(C_{1}\mu^{\theta}+C_{1}\eta^{-\theta})^{L}+\sqrt{L}m_{q}(r\rho/d)^{-2(q+1)}\Big{)}\|f\|_{H^{\theta}(\Omega)}.$
+Finally, we have the subsequent result on the cost of our multiscale algorithm, which is in line with a similar result for multiscale interpolation using hierarchical matrices, see [14].
+Theorem 6.3.
+Let $X_{1},\ldots,X_{L}$ be quasi-uniform and assume that the corresponding cardinalities $\{N_{\ell}\}_{\ell}$ form a geometric sequence, i.e., $N_{1}>0$ and $N_{\ell}=c_{\mathrm{geo}}^{\ell}N_{1}$ for some constant $c_{\mathrm{geo}}>1$.
+Then, the overall computational cost for assembling the linear system (21) and the number of non-vanishing matrix entries are both bounded by $\mathcal{O}(N_{L}\log^{2}N_{L})$, with the hidden constant depending on the dimension $m_{q}$ of $\mathcal{P}_{q}$.
+Proof.
+Fix a row index $\ell$.
+It has been shown within [1, Proof of Theorem 3.16] that the cost for the computation of $\mathbf{K}^{\Sigma,\rho}_{\ell,\ell^{\prime}}$ amounts to $\mathcal{O}\left(N_{\ell}\log N_{\ell^{\prime}}\right)$, which is also the number of non-vanishing entries of that row.
+Moreover, there are $\ell$ blocks per row.
+Hence, the cost for the computation of the $\ell$-th row of (21) is given by
+$\sum_{\ell^{\prime}=1}^{\ell}\mathcal{O}\left(N_{\ell}\log N_{\ell^{\prime}}\right)=\mathcal{O}\bigg{(}N_{\ell}\sum_{\ell^{\prime}=1}^{\ell}\log N_{\ell^{\prime}}\bigg{)}=\mathcal{O}\left(N_{\ell}\log^{2}N_{\ell}\right),$
+since $\ell=\mathcal{O}(\log N_{\ell})$ due to the geometric sequence property. Next, summing up the cost of the individual rows yields the overall cost
+$\sum_{\ell=1}^{L}\mathcal{O}\left(N_{\ell}\log^{2}N_{\ell}\right)=\mathcal{O}\bigg{(}\sum_{\ell=1}^{L}N_{\ell}\log^{2}N_{\ell}\bigg{)}=\mathcal{O}\bigg{(}\log^{2}N_{L}\sum_{\ell=1}^{L}N_{\ell}\bigg{)}=\mathcal{O}\left(N_{L}\log^{2}N_{L}\right),$
+again due to the geometric sequence property, i.e., $\sum_{\ell=1}^{L}N_{L}=\mathcal{O}(N_{L})$. This completes the proof. ∎
+To give an idea of the sparsity pattern of the samplet compressed system (21), Figure 1 visualizes the matrix pattern $\boldsymbol{K}^{\Sigma}_{\ell,\ell^{\prime}}$ using three levels of points $X_{1},X_{2},X_{3}$ arranged on a regular grid for the unit square, with cardinalities of 16 641, 66 049, and 263 169, respectively.
 Furthermore, we remark that, for a geometric sequence $\{N_{\ell}\}_{\ell}$, the cost for applying $\mathcal{T}$ and $\mathcal{T}^{\intercal}$, respectively, remains linear in $N_{L}$.
-Each block ${\boldsymbol{K}}^{\Sigma}_{\ell,\ell^{\prime}}$, $\ell\leq\ell^{\prime}$, can be computed with cost $\mathcal{O}(N_{\ell^{\prime}}+N_{\ell}\log N_{\ell})$ if $\{N_{\ell}\}_{\ell}$ forms a geometric sequence, cp.
-[1, Proof of Theorem 3.16].
-Moreover, there are $\mathcal{O}(\log N_{\ell^{\prime}})$ blocks per row, resulting in a computational cost of $\mathcal{O}(N_{\ell^{\prime}}\log N_{\ell^{\prime}})$ per row by the geometric sequence property of $\{N_{\ell}\}_{\ell}$.
-Summing up the cost for each row, exploiting the geometric sequence again, the overall computation cost for the assembly of the linear system (21) is bounded by $\mathcal{O}(N_{L}\log N_{L})$.
-To give an idea of the sparsity pattern of the samplet compressed system (21), Figure 1 visualizes the matrix pattern $\boldsymbol{K}^{\Sigma}_{\ell,\ell^{\prime}}$ using three levels of points $X_{1},X_{2},X_{3}$ arranged on a regular grid for the unit square, with cardinalities of 4 225, 66 049, and 1 050 625, respectively.
-Figure 1. Example of the sparsity patterns of the matrices $\boldsymbol{K}^{\Sigma}_{\ell,\ell^{\prime}}$ for ,$\ell,\ell^{\prime}\in\{1,2,3\}$ in the linear system (21).
+Figure 1. Example of the sparsity patterns of the matrices $\boldsymbol{K}^{\Sigma,\rho}_{\ell,\ell^{\prime}}$ for ,$\ell,\ell^{\prime}\in\{1,2,3\}$ of the compressed linear system from (21).
 7. Numerical results
 We present numerical examples of the proposed multiscale interpolation in samplets coordinates in both two and three dimensional settings.
-The parameters for the samplet compression are $\rho$, $q$, and $\kappa$.
-The parameter $\rho$ corresponds to the admissibility condition (18), which steers the a-priorily truncated matrix blocks.
+The parameters for the samplet compression are $\rho$ and $q$.
+The first one is associated with the admissibility condition (18), which determines the matrix blocks that are initially truncated.
 A larger $\rho$ results in a larger distance between clusters before they are admissible for compression.
-In our analysis, we adopt $\rho=d$, so that the bound in Equation (19) is exponentially governed by the number of vanishing moments.
-This number of vanishing moments corresponds to $q+1$, as specified in (15).
-The higher $q$, the more polynomial exactness is gained.
-Lastly, $\kappa$ serves as an a-posteriori compression threshold.
-After applying the samplet transform to the generalized Vandermonde matrix, we truncate entries smaller than $\kappa$, resulting in a sparser matrix representation.
+In our analysis, we adopt $\rho=d$, so that the bound in Equation (19) is exponentially governed by the number of vanishing moments, denoted as $q+1$, see (15).
+The higher $q+1$, the more polynomial exactness is gained.
+In the experiments, we vary also the parameter $\gamma$, defined as $\gamma=\eta\mu,$ where $\delta$ is the kernel rescaling factor, given by $\delta_{i}=\eta h_{X_{i}}$ and previously defined in (3), while $\eta$ represents the scaling factor for the fill distance between consecutive levels, i.e., $h_{X_{\ell+1}}\approx\mu h_{X_{\ell}}$.
 To measure the approximation quality, we consider the relative interpolation errors
-$\text{error}_{p,{\ell}}=\frac{\|f-s_{\ell}\|_{p}}{\|f\|_{p}}$
-for $p=2$ and $p=\infty$ at each of the levels $\ell=1,\ldots,L$. Furthermore, we compute the order of convergence, given by
-$\text{order}_{p,{\ell}}=\frac{\log{\left(\displaystyle\frac{\text{error}_{p,{\ell+1}}}{{\text{error}_{p,{\ell}}}}\right)}}{\log{\left(\displaystyle\frac{h_{\ell+1}}{h_{\ell}}\right)}}.$
+$\text{error}_{p,{m}}=\frac{\|f-\widehat{s}_{m}|X_{\text{eval}}\|_{\ell^{p}}}{\|f|X_{\text{eval}}\|_{\ell^{p}}}$
+for $p=2$ and $p=\infty$ at each of the levels $m=1,\ldots,L$ for a given set of evaluation points $X_{\text{eval}}$. Furthermore, we compute the order of convergence, given by
+$\text{order}_{p,{m}}=\frac{\log{\left(\displaystyle\frac{\text{error}_{p,{m+1}}}{{\text{error}_{p,{m}}}}\right)}}{\log{\left(\displaystyle\frac{h_{m+1}}{h_{m}}\right)}}.$
 We investigate three different scenarios in our numerical tests.
@@
 Finally, we extend to a three dimensional setting issuing from the Lucy surface model of the Stanford 3D Scanning Repository [^1] https://graphics.stanford.edu/data/3Dscanrep/.
-All computations have been performed at the Centro Svizzero di Calcolo Scientifico (CSCS) on a single node of the Alps cluster [^2] https://www.cscs.ch/computers/alps with two AMD EPYC 7742 @2.25GHz CPUs, up to 420GB of main memory and up to 20 cores.
+All computations have been performed at the Centro Svizzero di Calcolo Scientifico (CSCS) on a single node of the Alps cluster [^2] https://www.cscs.ch/computers/alps with two AMD EPYC 7742 @2.25GHz CPUs with 470GB of main memory.
+For the computations, we have used up to 16 cores.
 The implementation of the samplet compression is available as software package FMCA [^3] https://github.com/muchip/fmca.
@@
 on the unit square $\Omega=[0,1]^{2}$.
-To this end, we generate 10 sets of nested points within this domain, distributing them on a regular grid with size $h_{j}=2^{-j}$.
+To this end, we generate 10 sets of nested points within this domain, distributing them on a regular grid with size $h_{\ell}=2^{-\ell}$.
 The number of points per level and the corresponding fill-distances are provided in Table 1.
@@
 Table 1. Number of points per level and the corresponding fill-distances on $\Omega=[0,1]^{2}$ for the interpolation of Franke’s function.
-We have conducted computations using a-posteriori compression thresholds of $\kappa=10^{-4},10^{-6},10^{-8}$.
-For $\kappa=10^{-4}$, the error remains larger than $10^{-6}$, due to an increased compression error.
-Setting $\kappa=10^{-6}$, we have achieved an error of the order of $10^{-8}$.
-Further reducing the threshold to $\kappa=10^{-8}$ results in an error of $10^{-9}$.
-Even so, the latter results in a significantly worse compression rate of the generalized Vandermonde matrices.
-Therefore, we detail the results using $\kappa=10^{-6}$ for the threshold.
-We approximate the solution with the ${C}^{1}$-smooth Matérn-3/2 RBF and the ${C}^{2}$-smooth Matérn-5/2 RBF and different values of $\gamma$.
+The solution is approximated with the ${C}^{1}$-smooth Matérn-3/2 RBF and the ${C}^{2}$-smooth Matérn-5/2 RBF for different values of $\gamma$ and using $\rho=2$, $q=4$.
+Furthermore, we perform an a-posteriori thresholding of small matrix entries with a relative threshold of $10^{-8}$ with respect to the Frobenius norm, cp.
+Theorem 6.1.
 The error is evaluated at a fine grid with mesh-size $h=2^{-11}$.
@@
 Increasing the accuracy does not result in smaller errors but increases the number of iterations.
-Interpolation error, order and compression rate (%nz) as well as (non-preconditioned) conjugate gradient iterations (CG) are presented in the Tables 25.
+Interpolation error, order, and compression rate (measured as the percentage of nonzero entries, %nz), as well as (non-preconditioned) conjugate gradient iterations (CG), are presented in Tables 25.
 Matérn-3/2 RBF
@@
 1 $5.67\cdot 10^{-1}$ 55 5 $4.84\cdot 10^{-1}$ 55 6
-2 $1.38\cdot 10^{-1}$ 2.04 36 8 $5.29\cdot 10^{-2}$ 2.19 36 13
+2 $1.38\cdot 10^{-1}$ 2.04 40 8 $5.29\cdot 10^{-2}$ 2.19 40 13
 3 $3.16\cdot 10^{-2}$ 2.12 43 9 $9.20\cdot 10^{-3}$ 2.53 43 25
-4 $6.76\cdot 10^{-3}$ 2.23 47 10 $1.15\cdot 10^{-3}$ 3.00 47 34
-5 $1.48\cdot 10^{-3}$ 2.19 22 10 $2.51\cdot 10^{-4}$ 2.53 22 39
-6 $3.45\cdot 10^{-4}$ 2.10 7.8 10 $6.08\cdot 10^{-5}$ 2.05 7.6 39
-7 $8.30\cdot 10^{-5}$ 2.06 2.3 10 $1.30\cdot 10^{-5}$ 2.22 2.3 39
-8 $1.46\cdot 10^{-5}$ 2.51 0.62 10 $1.14\cdot 10^{-6}$ 3.51 0.62 39
-9 $3.14\cdot 10^{-6}$ 2.22 0.16 10 $1.04\cdot 10^{-7}$ 3.45 0.16 39
-10 $7.89\cdot 10^{-7}$ 1.99 0.04 10 $4.47\cdot 10^{-8}$ 1.21 0.04 39
-Table 2. Results for the interpolation of Franke’s function using the ${C}^{1}$-smooth Matérn-3/2 RBF and $\rho=2$, $q+1=5$, $\kappa=10^{-6}$ as samplet compression parameters.
+4 $6.76\cdot 10^{-3}$ 2.23 48 10 $1.15\cdot 10^{-3}$ 3.00 48 34
+5 $1.48\cdot 10^{-3}$ 2.19 22 10 $2.51\cdot 10^{-4}$ 2.53 28 39
+6 $3.45\cdot 10^{-4}$ 2.10 7.8 10 $6.08\cdot 10^{-5}$ 2.05 10 39
+7 $8.30\cdot 10^{-5}$ 2.06 2.3 10 $1.30\cdot 10^{-5}$ 2.22 3.2 39
+8 $1.46\cdot 10^{-5}$ 2.51 0.64 10 $1.14\cdot 10^{-6}$ 3.51 0.89 39
+9 $3.13\cdot 10^{-6}$ 2.22 0.17 10 $9.98\cdot 10^{-8}$ 3.52 0.23 39
+10 $7.85\cdot 10^{-7}$ 2.00 0.04 10 $3.47\cdot 10^{-8}$ 1.52 0.06 39
+Table 2. Results for the interpolation of Franke’s function using the ${C}^{1}$-smooth Matérn-3/2 RBF and $\rho=2$, $q=4$ as samplet compression parameters.
 Matérn-3/2 RBF
@@
 1 $4.84\cdot 10^{-1}$ 55 6 $4.88\cdot 10^{-1}$ 55 6
-2 $4.67\cdot 10^{-2}$ 3.37 36 17 $4.51\cdot 10^{-2}$ 3.43 36 20
-3 $7.97\cdot 10^{-3}$ 2.55 43 41 $7.72\cdot 10^{-3}$ 2.55 43 48
-4 $5.70\cdot 10^{-4}$ 3.81 47 64 $3.60\cdot 10^{-4}$ 4.43 47 97
-5 $1.09\cdot 10^{-4}$ 2.38 23 90 $5.75\cdot 10^{-5}$ 3.31 25 156
-6 $2.38\cdot 10^{-5}$ 2.19 8.2 96 $1.16\cdot 10^{-5}$ 2.31 8.9 187
-7 $4.57\cdot 10^{-6}$ 2.38 2.5 98 $2.07\cdot 10^{-6}$ 2.82 2.7 195
-8 $4.77\cdot 10^{-7}$ 3.26 0.68 99 $2.27\cdot 10^{-7}$ 3.19 0.74 197
-9 $5.97\cdot 10^{-8}$ 3.00 0.18 99 $7.85\cdot 10^{-8}$ 1.53 0.19 199
-10 $4.63\cdot 10^{-8}$ 0.37 0.05 99 $7.30\cdot 10^{-8}$ 0.10 0.05 200
-Table 3. Results for the interpolation of Franke’s function using the ${C}^{1}$-smooth Matérn-3/2 RBF and $\rho=2$, $q+1=5$, $\kappa=10^{-6}$ as samplet compression parameters.
+2 $4.67\cdot 10^{-2}$ 3.37 40 17 $4.51\cdot 10^{-2}$ 3.43 40 20
+3 $7.97\cdot 10^{-3}$ 2.55 43 40 $7.72\cdot 10^{-3}$ 2.55 43 48
+4 $5.70\cdot 10^{-4}$ 3.81 48 63 $3.60\cdot 10^{-4}$ 4.43 48 96
+5 $1.09\cdot 10^{-4}$ 2.38 35 90 $5.75\cdot 10^{-5}$ 3.31 37 155
+6 $2.38\cdot 10^{-5}$ 2.19 14 96 $1.16\cdot 10^{-5}$ 2.31 15 186
+7 $4.57\cdot 10^{-6}$ 2.38 4.4 98 $2.06\cdot 10^{-6}$ 2.49 5.0 195
+8 $4.74\cdot 10^{-7}$ 3.26 1.2 99 $2.16\cdot 10^{-7}$ 3.25 1.4 196
+9 $3.80\cdot 10^{-8}$ 3.64 0.32 98 $3.05\cdot 10^{-8}$ 2.83 0.38 197
+10 $1.33\cdot 10^{-8}$ 1.51 0.08 98 $1.41\cdot 10^{-8}$ 1.11 0.09 200
+Table 3. Results for the interpolation of Franke’s function using the ${C}^{1}$-smooth Matérn-3/2 RBF and $\rho=2$, $q=4$ as samplet compression parameters.
 Matérn-5/2 RBF
@@
 1 $5.45\cdot 10^{-1}$ 55 5 $4.72\cdot 10^{-1}$ 55 6
-2 $1.12\cdot 10^{-1}$ 2.28 36 8 $4.51\cdot 10^{-2}$ 1.39 36 16
-3 $2.27\cdot 10^{-2}$ 2.14 43 9 $7.73\cdot 10^{-3}$ 2.54 43 32
-4 $4.54\cdot 10^{-3}$ 2.32 47 10 $6.92\cdot 10^{-4}$ 3.48 47 42
-5 $1.00\cdot 10^{-3}$ 2.18 22 10 $1.49\cdot 10^{-4}$ 2.88 22 48
-6 $2.49\cdot 10^{-4}$ 2.01 7.6 10 $3.42\cdot 10^{-5}$ 2.12 7.7 48
-7 $6.13\cdot 10^{-5}$ 2.02 2.3 10 $6.88\cdot 10^{-6}$ 2.32 2.3 48
-8 $6.06\cdot 10^{-6}$ 3.34 0.62 9 $8.31\cdot 10^{-7}$ 3.05 0.62 48
-9 $1.15\cdot 10^{-6}$ 2.40 0.16 9 $6.15\cdot 10^{-8}$ 3.76 0.16 48
-10 $3.07\cdot 10^{-7}$ 1.91 0.04 9 $1.57\cdot 10^{-8}$ 1.97 0.04 48
-Table 4. Results for the interpolation of Franke’s function using the ${C}^{2}$-smooth Matérn-5/2 RBF and $\rho=2$, $q+1=5$, $\kappa=10^{-6}$ as samplet compression parameters.
+2 $1.12\cdot 10^{-1}$ 2.28 40 8 $4.51\cdot 10^{-2}$ 1.39 40 16
+3 $2.27\cdot 10^{-2}$ 2.14 43 9 $7.77\cdot 10^{-3}$ 2.54 43 32
+4 $4.54\cdot 10^{-3}$ 2.32 48 10 $6.92\cdot 10^{-4}$ 3.48 48 42
+5 $1.00\cdot 10^{-3}$ 2.18 22 10 $1.49\cdot 10^{-4}$ 2.21 25 48
+6 $2.49\cdot 10^{-4}$ 2.01 7.8 10 $3.42\cdot 10^{-5}$ 2.12 8.8 48
+7 $6.13\cdot 10^{-5}$ 2.02 2.3 10 $6.88\cdot 10^{-6}$ 2.31 2.7 48
+8 $6.06\cdot 10^{-6}$ 3.34 0.64 9 $8.31\cdot 10^{-7}$ 3.05 0.74 48
+9 $1.15\cdot 10^{-6}$ 2.40 0.17 9 $5.68\cdot 10^{-8}$ 3.87 0.19 48
+10 $3.07\cdot 10^{-7}$ 1.91 0.04 9 $6.86\cdot 10^{-9}$ 3.05 0.05 48
+Table 4. Results for the interpolation of Franke’s function using the ${C}^{2}$-smooth Matérn-5/2 RBF and $\rho=2$, $q=4$ as samplet compression parameters.
 Matérn-5/2 RBF
@@
 1 $4.79\cdot 10^{-1}$ 55 6 $9.16\cdot 10^{-1}$ 55 7
-2 $4.41\cdot 10^{-2}$ 3.44 36 20 $4.69\cdot 10^{-2}$ 4.29 36 25
-3 $7.13\cdot 10^{-3}$ 2.63 43 57 $7.47\cdot 10^{-3}$ 2.65 43 82
-4 $3.29\cdot 10^{-4}$ 4.43 47 97 $1.63\cdot 10^{-4}$ 2.19 47 172
-5 $6.14\cdot 10^{-5}$ 2.42 23 139 $3.96\cdot 10^{-5}$ 2.04 24 298
-6 $1.22\cdot 10^{-5}$ 2.33 8.0 151 $1.34\cdot 10^{-5}$ 1.56 8.8 370
-7 $2.11\cdot 10^{-6}$ 2.53 2.4 157 $2.06\cdot 10^{-6}$ 2.70 2.7 402
-8 $2.57\cdot 10^{-7}$ 3.03 0.70 159 $2.39\cdot 10^{-7}$ 3.10 0.73 416
-9 $4.06\cdot 10^{-8}$ 2.66 0.17 161 $7.56\cdot 10^{-8}$ 1.66 0.19 430
-10 $3.04\cdot 10^{-8}$ 0.42 0.04 162 $7.01\cdot 10^{-8}$ 0.11 0.05 430
-Table 5. Results for the interpolation of Franke’s function using the ${C}^{2}$-smooth Matérn-5/2 RBF and $\rho=2$, $q+1=5$, $\kappa=10^{-6}$ as samplet compression parameters.
+2 $4.41\cdot 10^{-2}$ 3.44 40 20 $4.69\cdot 10^{-2}$ 4.29 40 25
+3 $7.13\cdot 10^{-3}$ 2.63 43 58 $7.47\cdot 10^{-3}$ 2.65 43 83
+4 $3.29\cdot 10^{-4}$ 4.43 48 96 $1.63\cdot 10^{-4}$ 2.19 48 171
+5 $6.14\cdot 10^{-5}$ 2.42 35 139 $3.96\cdot 10^{-5}$ 2.04 37 301
+6 $1.22\cdot 10^{-5}$ 2.33 13 155 $1.34\cdot 10^{-5}$ 1.56 15 369
+7 $2.11\cdot 10^{-6}$ 2.53 4.2 157 $2.06\cdot 10^{-6}$ 2.70 4.8 403
+8 $2.57\cdot 10^{-7}$ 3.03 1.2 158 $2.39\cdot 10^{-7}$ 3.10 1.4 407
+9 $4.06\cdot 10^{-8}$ 2.66 0.31 159 $7.56\cdot 10^{-8}$ 1.66 0.36 436
+10 $3.04\cdot 10^{-8}$ 0.42 0.08 162 $7.01\cdot 10^{-8}$ 0.11 0.09 430
+Table 5. Results for the interpolation of Franke’s function using the ${C}^{2}$-smooth Matérn-5/2 RBF and $\rho=2$, $q=4$ as samplet compression parameters.
 The tables clearly demonstrate the expected exponential decay of the approximation error with increasing level for both error measures and both RBFs under consideration.
-Furthermore, we see the effect of a larger value of $\gamma$, which to a greater lengthscale parameter of the RBF, resulting in a larger condition number.
+Furthermore, we see the effect of a larger value of $\gamma$, which corresponds to a larger lengthscale parameter of the RBF, resulting in a larger condition number.
 For increasing $\gamma$, we require successively more iterations for the conjugate gradient method to achieve the prescribed accuracy, with up to 430 iterations for $\gamma=1$ and the Matérn-5/2 RBF.
-The deterioration of the convergence for $\gamma=0.75,1$ on the last level is caused by the samplet compression error, which is a directly to the chosen compression threshold $\kappa$.
-Figure 2 shows a visualization of the solution (left) and the residuals (right) at levels $\ell=1,4,7,10$ for $\gamma=1$ using the Matérn-3/2 RBF.
+To validate the theoretical bound on the computational cost of assembling the linear system stated in Theorem 2.1, Figure 2 shows the single threaded computation times obtained with $\gamma=1$ and the Matérn-3/2 RBF.
+The dashed lines correspond to the rates $N\log^{\alpha}N$ for $\alpha=1,2,3$.
+The results clearly confirm that the overall time scales as $N\log^{2}N$, in agreement with the theoretical analysis.
+[…]
+Figure 2. Comparison of the overall assembly time (blue) with theoretical growth rates $N\log^{\alpha}N$ for $\alpha=1,2,3$.
+Finally, Figure 3 shows a visualization of the solution (left) and the residuals (right) at levels $\ell=1,4,7,10$ for $\gamma=0.5$ using the Matérn-3/2 RBF.
 Refer to captionRefer to caption$\ell=1$Refer to captionRefer to caption$\ell=4$Refer to captionRefer to caption$\ell=7$Refer to captionRefer to caption$\ell=10$
@@
 We consider the L-shaped domain $\Omega=\left[-\frac{1}{2},\frac{1}{2}\right]^{2}\setminus\left(0,\frac{1}{2}\right]^{2}$ where we introduce polar coordinates $x=r\cos\phi$, $y=r\sin\phi$, with $r\geq 0$ and $\phi\in[\pi/2,2\pi]$.
-In this test, we aim at interpolating the function
-$u(r,\phi)=-r^{\frac{2}{3}}\sin\left(\frac{2\phi-\pi}{3}\right),$
-which solves the Laplace equation $\Delta u=0$.
-We select $h_{j}=2^{-j}$ as before.
+In this test, we aim at interpolating the function $u(r,\phi)=-r^{\frac{2}{3}}\sin\left(\frac{2\phi-\pi}{3}\right),$ which solves the Laplace equation $\Delta u=0$.
+We select $h_{\ell}=2^{-\ell}$ as before.
 Given the reduced smoothness of the function, we choose the ${C}^{0}$-smooth Matérn-1/2 RBF for its approximation.
-The error is evaluated at a fine grid with mesh-size $h=2^{-9}$ and the accuracy of the conjugate gradient method is set to $10^{-6}$.
-We have conducted experiments using $\kappa=10^{-4},10^{-6},10^{-8}$.
-While the magnitude of the errors remained consistent across these values, the compression rate in the final level increased as $\kappa$ decreased.
-Specifically for $\kappa=10^{-4}$, the compression rate is around $0.13\%$, for $\kappa=10^{-6}$, it is about $0.30\%$ and for $\kappa=10^{-8}$ it is approximately $0.43\%$.
-Consequently, we chose a higher threshold to reduce computational cost and storage.
-The number of points per level and the corresponding fill-distances are provided in Table 6.
+The error is evaluated on a fine grid with mesh size $h=2^{-9}$.
+The parameters for the samplet matrix compression and the accuracy of the conjugate gradient method are the same as in the previous example.
+Table 6 provides the number of points per level along with the corresponding fill distances.
+The results for $\gamma=0.5$ and $\gamma=1$ are summarized in Table 7.
 $\ell$ $N_{\ell}$ $h$
@@
 Table 6. Number of points per level and the corresponding fill-distances on the L-shaped domain for the interpolation of the solution to the Laplace equation.
-The results for $\gamma=0.5$ and $\gamma=1$ are summarized in Table 7.
 $\gamma=0.5$
 $\ell$ $\text{error}_{2}$ $\text{error}_{\infty}$ $\text{order}_{2}$ $\text{order}_{\infty}$nz CG
-1 $4.37\cdot 10^{-2}$ $5.98\cdot 10^{-2}$ 52 9
-2 $1.38\cdot 10^{-2}$ $3.17\cdot 10^{-2}$ 1.67 0.92 49 21
-3 $5.33\cdot 10^{-3}$ $2.04\cdot 10^{-2}$ 1.37 0.64 38 28
-4 $1.83\cdot 10^{-3}$ $1.29\cdot 10^{-2}$ 1.54 0.66 18 34
-5 $5.88\cdot 10^{-4}$ $8.15\cdot 10^{-3}$ 1.64 0.67 6.2 36
-6 $1.85\cdot 10^{-4}$ $5.04\cdot 10^{-3}$ 1.67 0.69 1.8 37
-7 $5.87\cdot 10^{-5}$ $3.11\cdot 10^{-3}$ 1.66 0.70 0.51 37
-8 $1.89\cdot 10^{-5}$ $1.61\cdot 10^{-3}$ 1.63 0.95 0.13 37
+1 $4.37\cdot 10^{-2}$ $5.98\cdot 10^{-2}$ 52 8
+2 $1.38\cdot 10^{-2}$ $3.17\cdot 10^{-2}$ 1.66 0.92 49 15
+3 $5.33\cdot 10^{-3}$ $2.04\cdot 10^{-2}$ 1.37 0.64 38 22
+4 $1.83\cdot 10^{-3}$ $1.29\cdot 10^{-2}$ 1.54 0.66 18 25
+5 $5.88\cdot 10^{-4}$ $8.15\cdot 10^{-3}$ 1.64 0.66 6.2 27
+6 $1.85\cdot 10^{-4}$ $5.06\cdot 10^{-3}$ 1.67 0.69 1.8 28
+7 $5.84\cdot 10^{-5}$ $3.17\cdot 10^{-3}$ 1.66 0.67 0.51 28
+8 $1.84\cdot 10^{-5}$ $1.59\cdot 10^{-3}$ 1.66 1.00 0.13 28
 $\gamma=1$
 $\ell$ $\text{error}_{2}$ $\text{error}_{\infty}$ $\text{order}_{2}$ $\text{order}_{\infty}$nz CG
-1 $3.39\cdot 10^{-2}$ $6.06\cdot 10^{-2}$ 48 13
-2 $1.27\cdot 10^{-2}$ $3.94\cdot 10^{-2}$ 1.41 0.62 49 28
-3 $4.15\cdot 10^{-3}$ $2.50\cdot 10^{-2}$ 1.62 0.66 37 49
-4 $1.29\cdot 10^{-3}$ $1.58\cdot 10^{-2}$ 1.68 0.66 18 73
-5 $3.97\cdot 10^{-4}$ $9.94\cdot 10^{-3}$ 1.70 0.67 6.0 93
-6 $1.23\cdot 10^{-4}$ $6.13\cdot 10^{-3}$ 1.70 0.70 1.8 102
-7 $3.83\cdot 10^{-5}$ $3.72\cdot 10^{-3}$ 1.68 0.72 0.12 104
-8 $1.20\cdot 10^{-5}$ $1.95\cdot 10^{-3}$ 1.68 0.76 0.09 105
-Table 7.
-Results for the interpolation of the solution of the Laplace equation on the L-shaped domain using the ${C}^{0}$-smooth Matérn-1/2 RBF and $\rho=2$, $q+1=5$, $\kappa=10^{-4}$ as samplet compression parameters.
+1 $3.39\cdot 10^{-2}$ $4.81\cdot 10^{-2}$ 52 9
+2 $1.27\cdot 10^{-2}$ $3.12\cdot 10^{-2}$ 1.42 0.62 51 21
+3 $4.15\cdot 10^{-3}$ $1.98\cdot 10^{-2}$ 1.61 0.66 48 37
+4 $1.29\cdot 10^{-3}$ $1.25\cdot 10^{-2}$ 1.69 0.66 39 54
+5 $3.97\cdot 10^{-4}$ $7.90\cdot 10^{-3}$ 1.70 0.66 19 69
+6 $1.22\cdot 10^{-4}$ $4.90\cdot 10^{-3}$ 1.70 0.69 6.7 77
+7 $3.79\cdot 10^{-5}$ $3.04\cdot 10^{-3}$ 1.69 0.69 1.9 79
+8 $1.13\cdot 10^{-5}$ $1.15\cdot 10^{-3}$ 1.75 1.40 0.44 79
+Table 7. Results for the interpolation of the solution of the Laplace equation on the L-shaped domain using the ${C}^{0}$-smooth Matérn-1/2 RBF and $\rho=2$, $q=4$ as samplet compression parameters.
 The compression rates are similar to the previous example, also in case of the less smooth Matérn-1/2 RBF.
 The approximation error and the corresponding order are reduced, however.
-The number of conjugate gradient iterations stays rather low with a maximum of 37 iterations for $\gamma=0.5$ and of 105 iterations for $\gamma=1$.
+The number of conjugate gradient iterations stays rather low with a maximum of 28 iterations for $\gamma=0.5$ and of 79 iterations for $\gamma=1$.
 This can be explained by the slower decay of the Matérn-1/2 kernel’s eigenvalues, resulting in less ill-conditioned generalized Vandermonde matrices.
@@
 Refer to captionRefer to caption$\ell=2$Refer to captionRefer to caption$\ell=4$Refer to captionRefer to caption$\ell=6$Refer to captionRefer to caption$\ell=8$
-Figure 3.
-Solution (left) and corresponding residuals (right) for the interpolation of the solution to the Laplace equation on the L-shaped domain for levels $\ell=2,4,6,8$ using $\gamma=1$ and $\kappa=10^{-4}$.
+Figure 4. Solution (left) and corresponding residuals (right) for the interpolation of the solution to the Laplace equation on the L-shaped domain for levels $\ell=2,4,6,8$.
 7.3. Function on a point cloud in three spatial dimensions
@@
 Unlike previous tests, this experiment is more general, as the points are not arranged on a uniform grid, they are not necessarily nested, and they are not necessarily quasi-uniform.
-As test function, we e interpolate
-$f({\boldsymbol{x}})=\frac{1}{\sqrt[4]{\|{\boldsymbol{x}}-{\boldsymbol{x}}_{0}\|_{2}}+10^{-4}}$
-The point ${\boldsymbol{x}}_{0}$ is selected near the torch that Lucy holds in her left hand.
-To optimize performance, we use a relatively small value for the parameter $\gamma$, reduce the number of vanishing moments, and set a higher a-posteriori threshold.
+As test function, we interpolate $f({\boldsymbol{x}})=\big{(}\sqrt[4]{\|{\boldsymbol{x}}-{\boldsymbol{x}}_{0}\|_{2}}+10^{-4}\big{)}^{-1}$ The point ${\boldsymbol{x}}_{0}$ is selected near the torch that Lucy holds in her left hand.
+To optimize performance, we use a relatively small value for the parameter $\gamma$ and set $\rho=3$, $q=2$, while the a-posteriori threshold is set to $10^{-5}$.
 These adjustments help decrease both computational time and the number of non-zero entries.
@@
 The number of points per level and the corresponding fill-distances are provided in Table 8.
-$\ell$ $N_{\ell}$ $h$ (nested) $h$ (not-nested)
+$\ell$ $N_{\ell}$ $h$ (nested) $h$ (non-nested)
 1 20 0.25 0.34
@@
 The evaluation set remains fixed throughout all experiments to ensure consistency in error measurement.
-The results using the Matérn-1/2 and Matérn-3/2 RBFs are shown in Tables 9 and 10.
+The results using the Matérn-1/2 and Matérn-3/2 RBFs are shown in Tables 9 and 10, respectively.
 $\gamma=0.2$
-nested not-nested
+nested non-nested
 $\ell$ $\text{error}_{2}$ $\text{order}_{2}$nz CG $\text{error}_{2}$ $\text{order}_{2}$nz CG
@@
 5 $7.94\cdot 10^{-4}$ 1.56 0.03 57 $2.51\cdot 10^{-4}$ 1.74 0.03 69
-Table 9. Results for the interpolation of a test function on the Lucy point cloud for the $C^{0}$-smooth Matérn-1/2 RBF using $\rho=3$, $q+1=3$, $\kappa=10^{-4}$ as samplet compression parameters.
+Table 9. Results for the interpolation of a test function on the Lucy point cloud for the $C^{0}$-smooth Matérn-1/2 RBF using $\rho=3$, $q=2$ as samplet compression parameters.
 $\gamma=0.2$
-nested not-nested
+nested non-nested
 $\ell$ $\text{error}_{2}$ $\text{order}_{2}$nz CG $\text{error}_{2}$ $\text{order}_{2}$nz CG
@@
 5 $8.57\cdot 10^{-4}$ 1.53 0.03 349 $1.61\cdot 10^{-4}$ 1.64 0.03 430
-Table 10. Results for the interpolation of a test function on the Lucy point cloud for the $C^{1}$-smooth Matérn-3/2 RBF using $\rho=3$, $q+1=3$, $\kappa=10^{-4}$ as samplet compression parameters.
+Table 10. Results for the interpolation of a test function on the Lucy point cloud for the $C^{1}$-smooth Matérn-3/2 RBF using $\rho=3$, $q=2$ as samplet compression parameters.
 We observe that the $C^{0}$-smooth Matérn-1/2 RBF performs better in this scenario, primarily due to the improved conditioning of the associated generalized Vandermonde matrix.
-In both cases, the approximation error reaches the compression error of $10^{-4}$.
-A visualization of the solution on all five levels $\ell=1,\ldots,5$ using the Matérn-1/2 RBF and not-nested points is depicted in Figure 4.
+In both cases, the approximation error is higher than the compression error $\kappa=10^{-5}$.
+A visualization of the solution on all five levels $\ell=1,\ldots,5$ using the Matérn-1/2 RBF and non-nested points is depicted in Figure 5.
 Refer to captionRefer to caption$\ell=1$$\ell=2$Refer to captionRefer to caption$\ell=3$$\ell=4$Refer to caption$\ell=5$
@@
 In particular, we have extended existing results on the boundedness of the condition numbers of the generalized Vandermonde matrices within the multiscale approximation towards globally supported RBFs.
-For appropriately chosen lengthscale parameters, the condition numbers stay independent of the level, which allows for an efficient solution of the resulting linear system by the conjugate gradient method with a bounded number of iterations for a given accuracy.
-The resulting algorithm has an overall cost of $\mathcal{O}(N\log N)$ for $N$ quasi-uniform data sites, given that the points per level form a geometric sequence.
-The numerical results demonstrate that the presented approach achieves a similar accuracy to the multiscale approach by compactly supported RBFs, while overcoming previous limitations that made the use of globally supported RBFs impractical for large-scale problems.
+For appropriately chosen lengthscale parameters, the condition numbers of the diagonal blocks of the multiscale system stay bounded independently of the level.
+As a consequence, the diagonally scaled multiscale system becomes well conditioned.
+We have exploit this fact and derived a general error estimate bounding the consistency error issuing from a numerical approximation of the multiscale system.
+This result particularly allows to rigorously bound the consistency error arising from the samplet matrix compression.
+The sparse representation of the multiscale system in samplet coordinates allows for an efficient solution of the linear system by the conjugate gradient method with a bounded number of iterations for a given accuracy, due to the well conditioning.
+The resulting algorithm has an overall cost of $\mathcal{O}(N\log^{2}N)$ for $N$ quasi-uniform data sites, given local approximation spaces with exponentially decreasing dimension with decreasing level.
+The numerical results demonstrate that the presented approach achieves a similar accuracy to the multiscale approach using compactly supported RBFs, as presented in [29], where the corresponding experiments can be found.
+Meanwhile, our approach overcomes previous limitations that made the use of globally supported RBFs impractical for large-scale problems.
 We have demonstrated that the present multiscale approach can handle datasets with millions of data in both two and three spatial dimensions.
@@
 In all cases, we achieved accurate results that would have been computationally infeasible without the samplet matrix compression.
-Acknowledgement. SA and MM were funded by the SNSF starting grant “Multiresolution methods for unstructured data” (TMSGI2_211684).